### v1

# Local Dissipativity Analysis of Nonlinear Systems

Abstract
[…]
keywords:
robust control, nonlinear systems, dissipativity
thanks: This work was supported by the ONR YIP Grant No. N00014-23-1-2043 and NSF Grant No. 2303158.
and
1 Introduction
[…]
io stability theory is an important tool in robust control. It uses a compositional approach to stability analysis – that is, the closed loop IO stability of a network depends on the coarse, open loop IO characterizations of each individual subsystem [moylan1978stability, vidyasagar1981input]. This enables flexibility when addressing different types of subsystems – linear, nonlinear, time-varying, etc – with the additional benefit that IO characterizations can be selected that are robust to various sources of uncertainty. But therein lies the challenge: how do we find a valid IO characterization? For nonlinear systems, determining the IO characterization is an involved and, often, model specific process. For nonlinear, control affine systems, this work provides a comprehensive method to determine one such IO characterization: local QSR dissipativity, which encompasses local versions of a range of well-known IO characterizations, including passivity [brogliato2007dissipative], $\mathcal{L}_{2}$ gain [van19922], conic sectors [zames1966input], and more [vidyasagar1977l2].
For linear systems, we can establish IO properties through frequency-domain analysis [brogliato2007dissipative], solving LMIs that arise from Hamilton Jacobi Inequalitys [Gupta1996], or data-driven analysis [koch2021provably]. While IO theory takes an operator theoretic approach to dynamical systems analysis [zames1966input, vidyasagar1981input, moylan1978stability], the alternative state-space perspective on QSR dissipativity from [willems1972dissipative, hill1976stability] provides an avenue for analysis by linking a systems internal and supplied energy through a so-called storage function. For linear systems, this relationship led to the celebrated Kalman-Yakubovich-Popov (KYP) Lemma [brogliato2007dissipative], which provides necessary and sufficient conditions for dissipativity. The KYP Lemma equates dissipativity with the existence of a positive, semi-definite, quadratic storage function, which can be found by solving LMIs via optimization. Analogous to linear systems, dissipativity of a nonlinear, control affine system is equivalent to the existence of a positive semi-definite storage function [hill1976stability, hill1980connections]. However, because the storage function lacks a set functional form, there is no systematic way to find it, so model-dependent ad-hoc ingenuity is needed. As a result, much effort has been focused on developing model agnostic analysis methods for nonlinear systems.
IO analysis of nonlinear systems requires synthesis of the system’s storage function that adheres to the associated HJI. The problem of IO analysis of nonlinear systems can be approached from an optimal control perspective – using dynamic programming (DP) schemes to approximate a system’s storage function [james1995numerical, dower2008dynamic]. However, [dower2008dynamic] is limited to determining gain over a finite horizon, while [james1995numerical] uses iterative methods that require a good initialization and do not always have convergence guarantees. Another approach is to transform the nonlinear system into an easier form for analysis [sutavani2023small, zahedzadeh2009upper]. For example, in [sutavani2023small], nonlinear systems are approximated using Koopman operators so that analysis tools for linear systems can be used [sutavani2023small]. These methods are only valid in a limited region of the state space where the transformation is valid, and often require ad hoc ingenuity to determine the transformation.
[…]
Recent work formulated a nonconvex optimization problem that bounds the gain a of a nonlinear system by searching for a CPA storage function that satisfies the HJI [lavaei2022simultaneous] Constraints were imposed on the storage function at each vertex of the triangulation to ensure certain properties, leveraging a similar process as in previous work on Lyapunov function synthesis [giesl2012construction, giesl2014revised]. Specifically, the optimization problem in [lavaei2022simultaneous] depends on a specific error term that, when imposed with the HJI on each vertex of the triangulation, ensures the HJI holds for the entire region. However, the HJI and its error term are polynomial in design variables – leading to a conservative, nonconvex optimization problem that requires methods like IO to solve. Moreover, specific requirements on the error bound restricted the optimization problem to a limited class of nonlinear systems – notably precluding linear and control affine terms.
Contributions
This paper determines the local QSR dissipativity characterization of nonlinear, control affine systems via synthesis of a CPA storage function that adheres to the HJI over a region of the state space. This analysis is accomplished via a convex optimization problem that simultaneously searches for the QSR parameters and CPA storage function that satisfies that the HJI – ensuring that the inequality holds for the entirety of the region. In essence, our approach triangulates a region of the state space, and then upper bounds the nonlinear dynamics of the system on each simplex with a linear system. We are then able to synthesize a storage function across the linearized systems.
In Section 3, we develop a novel LMI error bound to impose LMI constraints on a CPA function on a triangulation. We also develop a method to enforce inequalities on a closed norm ball about the origin, which allows us to consider quadratic storage functions. The development of these error bounds builds on work in [strong2024improved]. In Section 4, we develop conditions for a CPA storage function to adhere to the HJI. Because of the LMI error bounds, we are able to reformulate the HJI as an LMI– convexifying the previous optimization problem. Optimization problems are formulated to search for general dissipativity characterizations of nonlinear systems. In Section 5 theoretical results are developed to confirm that if a nonlinear control affine system is strictly locally dissipative, then this optimization problem can be used to verify dissipativity characteristics for either CPA or quadratic storage functions. Finally, numerical results are given in Section 6.
Notation
The interior, boundary, and closure of the set $\Omega\subset\mathbb{R}^{n}$ are denoted as $\Omega^{\circ}$, $\delta\Omega$, and $\bar{\Omega}$. The symbol $\mathfrak{R}^{n}$ denotes the set of all compact subsets $\Omega\subset\mathbb{R}^{n}$ satisfying i) $\Omega^{\circ}$ is connected and contains the origin and ii) $\Omega=\bar{\Omega}^{\circ}.$ Let $\mathbb{R}_{+}$ represent the set of non-negative real numbers. The notation $\mathbb{Z}_{a}^{b}$ ($\mathbb{Z}_{\bar{a}}^{\underline{b}}$) denote the set of integers between $a$ and $b$ inclusive (exclusive). The p-norm of a vector $x\in\mathbb{R}^{n}$ is denoted as $\left\lVert\cdot\right\rVert_{p},$ where $p\in\mathbb{Z}_{1}^{\infty}.$ Let $B_{\epsilon}(0)$ denote a closed two-norm ball about the origin with radius $\epsilon.$
The positive definiteness (semi-definiteness) of a matrix is denoted by $\mathbf{P}\succ 0$ ($\mathbf{P}\succeq 0$). Identity and zero matrices are denoted as $\mathbf{I}$ and $\mathbf{0}.$ Let $1_{n}$ denote a vector of ones in $\mathbb{R}^{n},$ and let $\text{He}(\mathbf{X})=\mathbf{X}+\mathbf{X}^{\top}.$ Recall that for a matrix operator norm, $\left\lVert\mathbf{A}\mathbf{B}\right\rVert\leq\left\lVert\mathbf{A}\right\rVert\left\lVert\mathbf{B}\right\rVert$ [petersen2008matrix]. From Lemma 2.3 of [giesl2012construction], if $\bar{h}$ is the maximum absolute value of any element within the matrix $\mathbf{H}\in\mathbb{R}^{n\times m}$, then $\left\lVert\mathbf{H}\right\rVert_{2}\leq\sqrt{nm}\bar{h}.$
Let the $\mathcal{L}_{2}^{n}$ space be a set containing all measurable functions $\mathbf{u}:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ that satisfy $\int_{0}^{\infty}\mathbf{u}(t)^{\top}\mathbf{u}(t)dt<\infty$. The extended $\mathcal{L}^{n}_{2e}$ space is defined as the set of all measurable functions $\mathbf{u}:\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$ for which the truncation to $t\in[0,T]$ is in $\mathcal{L}_{2}^{n}$ for all $T\geq 0$.
For the function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n},$ let $f^{(k)}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ denote the $k^{th}$ element of $f,$ and let $\mathbf{H}_{f^{(k)}}$ denote the Hessian for this element, where $k\in\mathbb{Z}_{1}^{n}.$ Further, let $f{\in}\mathcal{C}^{r}(\mathcal{X})$ denote a real valued function is $r$-times differentiable over the domain $\mathcal{X}$. The right-hand (left-hand) upper Dini-derivatives for a function, $f$, is defined $D^{+}f(\mathbf{y})\vcentcolon=\lim\sup_{k\rightarrow 0^{+}}\frac{f(\mathbf{x}{+}kg(\mathbf{x})){-}f(\mathbf{x})}{k}$
$\left(\lim\sup_{k\rightarrow 0^{-}}\frac{f(\mathbf{x}{+}kg(\mathbf{x})){-}f(\mathbf{x})}{k}\right),$ where $k{\in}\mathbb{R}$ and $\dot{\mathbf{x}}{=}g(\mathbf{x})$ [giesl2014revised].
2 Problem Statement and Background
This paper aims to characterize the local dissipativity properties of the mapping $\mathcal{G}(x_{0}):\mathcal{L}_{2e}^{m}\rightarrow\mathcal{L}_{2e}^{p}$, where $\mathbf{y}=\mathcal{G}(x_{0})\mathbf{u}$ is defined by $x(0)=x_{0}\in\mathbb{R}^{n}$ and
$\mathcal{G}(x_{0}):\begin{cases}\dot{x}&=f(x)+\bar{G}(x)u,\\ y&=h(x)+\bar{J}(x)u.\end{cases}$ (1)
Here, $\bar{G}(\mathbf{x})=\mathbf{B}+G(\mathbf{x})$ and $\bar{J}(\mathbf{x})=\mathbf{D}+J(\mathbf{x}).$ Let $f(0)=0,$ $G(0)=0,$ $h(0)=0,$ and $J(0)=0$.
We will establish local dissipativity by developing constraints on CPA storage functions implying QSR dissipativity. Local dissipativity will be sought for a family of mappings, $\mathcal{G}_{\Omega}\vcentcolon=\{\mathcal{G}(x_{0})\mid x_{0}\in\Omega\},$ defined over a robustly positive invariant subset of the state space, $\Omega\subset\mathbb{R}^{n}$, for inputs in a non-empty set $\mathcal{U}_{\Omega}\subseteq\mathcal{L}_{2}^{m}$, where $\mathcal{U}_{\Omega}\vcentcolon=\{\mathcal{U}_{\Omega}(x_{0})\mid x_{0}\in\Omega\}.$ Here, if $x_{0}\in\Omega$ and $\mathbf{u}\in\mathcal{U}_{\Omega}(x_{0})$, then $x(t)\in\Omega$ for all $t\geq 0$.
The following sections detail the necessary definitions and tools for local dissipativity analysis.
2.1 Local Dissipativity
Hill and Moylan informally defined a dissipative system as a “system that absorbs more energy from the external world than it supplies” [hill1976stability]. Essentially, any change in the system’s internal energy (represented by the positive semi-definite “storage function”) is bounded above by a function of the input signal to the system and the resulting output, as shown below.
Definition 1.
[willems1972dissipative, Definition 2]: A dynamical system (1) with a locally integrable supply rate $w(t)=w(u,y)$ is dissipative, if there exists a non-negative function, $V:\mathbb{R}^{n}\rightarrow\mathbb{R}_{+}$, called the storage function, such that for all $(t_{0},t_{1})\in\mathbb{R}_{+}$, $x\in\mathbb{R}^{n}$, and $\mathbf{u}\in\mathcal{L}_{2}^{m}$
$V(x_{0})+\int_{t_{0}}^{t_{1}}w(t)dt\geq V(x_{1}),$ (2)
where $x_{1}$ and $y(t)$ are the result of applying (1) from time $t_{0}$ to $t_{1}$ with input $\mathbf{u}(t).$
While dissipativity was originally defined globally, there are many nonlinear systems for which dissipativity is a local property about an attractor [hill1980connections, vidyasagar1982new]. Different definitions of local dissipativity have been used throughout the literature with various input signals, attractors, and system definitions [hill1980connections, vidyasagar1982new, pota1993stability]. Below, Definition 2 gives a local QSR dissipativity that brings together several concepts in the literature [hill1980connections, pota1993stability].
Definition 2.
The mapping (1) is locally QSR dissipative on $\Omega\in\mathfrak{R}^{n}$, if the following conditions are satisfied:
  1. 1.
    The set $\mathcal{U}_{\Omega}(x_{0})$ is non-empty, full dimensional, and contains a neighborhood of the origin.
  2. 2.
    Given the supply rate function
    $w(u,y)=y^{\top}\mathbf{Q}y+2y^{\top}\mathbf{S}u+u^{\top}\mathbf{R}u,$ (3)
    the inequality
    $\int_{0}^{T}w(u,y)\geq\beta$ (4)
    is satisfied for all $\mathbf{u}\in\mathcal{U}_{\Omega}(x_{0})$ and $T\geq 0$, where $\beta\in\mathbb{R}$, $\mathbf{Q}=\mathbf{Q}^{\top}\in\mathbb{R}^{p\times p}$, $\mathbf{S}\in\mathbb{R}^{p\times m}$, and $\mathbf{R}=\mathbf{R}^{\top}\in\mathbb{R}^{m\times m}.$
The family of mappings $\mathcal{G}_{\Omega}$ is locally dissipative if all of its elements are.
The quadratic supply rate in (3) admits sufficient conditions for QSR dissipativity of smooth, nonlinear, control affine systems [hill1976stability, Theorem 1], [van19922, Theorem 2].However, these were originally formulated for global dissipativity. Local dissipativity can be established instead by applying the proof strategy from [van19922, Theorem 2] while restricting the initial state to a set, $x_{0}\in\Omega$, that is robustly positive invariant to inputs in $u\in\mathcal{U}_{\Omega}\neq\emptyset$ and using a more general storage function. The resulting conditions implying local dissipativity of (1) in $\Omega$ when $\mathcal{U}_{\Omega}$ are summarized next, providing an LMI that mirrors the KYP Lemma.
Theorem 3.
[hill1976stability, Theorem 1], [van19922, Theorem 2] Consider (1) on $\Omega\subset\mathbb{R}^{n}$, where $\mathcal{U}_{\Omega}(x_{0})$ is non-empty for all $x_{0}\in\Omega$. If there exists a non-negative, smooth function, $V:\Omega\rightarrow\mathbb{R}_{+},$ that satisfies
$\begin{bmatrix}\nabla V^{\top}f(x)&\frac{1}{2}\nabla V^{\top}\bar{G}(x)-h(x)^{\top}\mathbf{S}&h(x)^{\top}\\ *&\hat{\mathbf{R}}&\bar{J}(x)^{\top}\\ *&*&\mathbf{Q}^{-1}\end{bmatrix}\preceq 0,$ (5)
where $\hat{\mathbf{R}}=-\mathbf{R}-\mathbf{S}^{\top}\bar{J}(x)-\bar{J}(x)^{\top}\mathbf{S}$, then $\mathcal{G}_{\Omega}$ is locally dissipative on $\Omega$ with respect to (3).
{pf}
Let $x_{0}\in\Omega$ and $\mathbf{u}\in\mathcal{U}_{\Omega},$ then [hill1976stability, Equation 6] holds where $V(x)=\phi(x).$ Thus, $\mathcal{G}_{\Omega}$ adheres to [hill1976stability, Theorem 1] for $x_{0}\in\Omega$ and $\mathbf{u}\in\mathcal{U}.$
2.2 Continuous Piecewise Affine Functions
The storage function $V$ is non-unique and has no set functional form [willems1972dissipative]. Further, the QSR dissipativity condition is a function of $x\in\Omega$. CPA functions will help us in the nontrivial task of searching for a storage function which satisfies (5) across $\Omega$. Inspired by the CPA Lyapunov function search techniques of [giesl2012construction, giesl2014revised], we will use this flexible functional form for candidate storage functions.
Definition 4.
[giesl2014revised] Affine independence: A collection of $m$ vectors $\{x_{0},x_{1},\ldots,x_{m}\}\subset\mathbb{R}^{n}$ is affinely independent if $x_{1}-x_{0},\ldots,x_{m}-x_{0}$ are linearly independent.
Definition 5.
[giesl2014revised] $n$ - simplex: A simplex, $\sigma$, is defined the convex hull of $n+1$ affinely independent vectors, $co\{x_{j}\}_{j=0}^{n}$, where each vector, $x_{j}\in\mathbb{R}^{n}$, is a vertex.
Definition 6.
[giesl2014revised] Triangulation: Let $\mathcal{T}{=}\{\sigma_{i}\}_{i=1}^{m_{\mathcal{T}}}{\in}\mathfrak{R}^{n}$ represent a finite collection of $m_{\mathcal{T}}$ simplexes, where the intersection of any two simplexes is a face or an empty set.
Let $\mathcal{T}=\{\sigma_{i}\}_{i=1}^{m_{\mathcal{T}}}.$ Further, let $\{x_{i,j}\}_{j=0}^{n}$ be a simplex’s ($\sigma_{i}$) vertices. The choice of $x_{i,0}$ in $\sigma_{i}$ is arbitrary unless $0\in\sigma_{i},$ in which case $x_{i,0}=0$ [giesl2014revised]. The vertices of the triangulation $\mathcal{T}$ in a set $\Omega$ are denoted as $\mathbb{E}_{\Omega}.$ Let $\mathcal{T}_{0}$ denote the simplexes in $\mathcal{T}$ containing 0 and $\mathcal{T}_{\Omega\setminus\{0\}}$ denotes simplices that do not contain 0.
An important characteristic of a CPA function is that it is defined uniquely over all of $\Omega\in\mathfrak{R}^{n}$ by its values on the finite set of vertices in $\mathcal{T},$ as detailed in Lemma 7.
Lemma 7.
[giesl2014revised, Remark 9] Consider the triangulation $\mathcal{T}=\{\sigma_{i}\}_{i=1}^{m_{\mathcal{T}}},$ where $\sigma_{i}=\text{co}(\{x_{i,j}\}_{j=0}^{n})$, and a set $\mathbf{W}=\{W_{x}\}_{\mathbf{x}\in\mathbb{E_{\mathcal{T}}}}\subset\mathbb{R},$ where $W(x)=W_{x},\forall x\in\mathbb{E}_{\mathcal{T}}.$ For simplex $\sigma_{i}$, let $\mathbf{X}_{i}\in\mathbb{R}^{n\times n}$ be a matrix that has $x_{i,j}-x_{i,0}$ as its $j$-th row and $\bar{W}_{i}\in\mathbb{R}^{n}$ be a vector that has $W_{x_{i,j}}-W_{x_{i,0}},$ as its $j$-th element. The function $W(x)=x^{\top}\mathbf{X}_{i}^{-1}\bar{W}_{i},$ is the unique CPA interpolation of $\mathbf{W}$ on $\mathcal{T}$ for $x\in\sigma_{i}$.
Like in [giesl2012construction], Taylor’s theorem (Theorem 8) is used to bound the change between a function $\zeta\in\mathcal{C}^{2}$ when evaluated at an arbitrary point versus the vertices on a simplex.
Theorem 8 (Taylor’s Theorem[fitzpatrick2009advanced]).
Consider $\zeta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, where $\zeta\in\mathcal{C}^{2}.$ For any $x,y\in\mathbb{R}^{n},$
$\begin{split}\zeta^{(k)}(x)=&\zeta^{(k)}(y){+}\langle\nabla\zeta^{(k)}(x),x-y\rangle{+}\frac{1}{2}\overline{\mathbf{H}}_{\zeta^{(k)}}(x,y),\end{split}$ (6)
where $\overline{\mathbf{H}}_{\zeta^{(k)}}(x,y)=\langle\mathbf{H}_{\zeta^{(k)}}(z)(x-y),(x-y)\rangle$ and $z$ is some convex combination of $x$ and $y.$
Let $\overline{\mathbf{H}}_{\zeta}(x,y)=\begin{bmatrix}\overline{\mathbf{H}}_{\zeta^{(1)}}(x,y)&\ldots&\overline{\mathbf{H}}_{\zeta^{(n)}}(x,y)\end{bmatrix}^{\top}$ be known as the Taylor remainder.
Although storage functions must satisfy more complicated constraints than Lyapunov functions, the error bounds used to establish CPA Lyapunov functions in [giesl2012construction, giesl2014revised] be useful here.
Lemma 9.
[giesl2012construction, Proposition 2.2 and Lemma 2.3] Consider $\Omega\in\mathfrak{R}^{n}$. Let $\zeta{:}\Omega{\rightarrow}\mathbb{R}^{n}$ where $\zeta\in\mathcal{C}^{2}(\Omega)$. Then, for any $x\in\sigma=\text{co}(\{x_{j}\}_{j=0}^{n}){\in\Omega}$,
$\left\lVert\zeta(x){-}\sum_{j=0}^{n}\lambda_{j}\zeta(x_{j})\right\rVert_{\infty}\!\!\!\!\!{=}$ $\frac{1}{2}\!\sum_{j=0}^{n}\!\lambda_{j}\!\left\lVert\overline{\mathbf{H}}_{\zeta}(x,x_{0}){-}\overline{\mathbf{H}}_{\zeta}(x_{j},x_{0})\right\rVert_{\infty}$
$\!\!\!\!\!\leq$ $\frac{1}{2}n\beta\sum_{j=0}^{n}\lambda_{j}\bar{c}_{j},$ (7)
where $\{\lambda_{j}\}_{j=0}^{n}\in\mathbb{R}_{0}^{1}$ is the set of unique coefficients satisfying $x=\sum_{j=0}^{n}\lambda_{j}x_{j}$ with $\sum_{j=0}^{n}\lambda_{j}=1$,
$\beta\geq\max_{p,q,r\in\mathbb{Z}_{1}^{n}}\max_{\xi\in\sigma}\left\lvert\frac{\partial^{2}\zeta^{(p)}(x)}{\partial x^{(q)}\partial x^{(r)}}\Bigr|_{x=\xi}\right\rvert,\text{ and }$ (8)
$\bar{c}_{j}{=}\left\lVert x_{j}{-}x_{0}\right\rVert\!(\max_{k\in\mathbb{Z}_{1}^{n}}\left\lVert x_{k}{-}x_{0}\right\rVert\!{+}\!\left\lVert x_{j}{-}x_{0}\right\rVert).$ (9)
In Section A, Lemma 17 derives a similar, but tighter, error bound than the one presented above. The proof also uses Taylor’s theorem to bound the change in a function across a simplex. But while Lemma 9 performs a Taylor expansion about a vertex point of the simplex, $x_{0}$, Lemma 17 (Section A) performs the expansion about an arbitrary point within the simplex. Importantly, the behavior of these error bounds differs at the origin. When $x_{j}=0$, $c_{j}=0$ in Lemma 9, while $c_{j}\neq 0$ in Lemma 17. Thus, we will use the tighter bound from Lemma 17 throughout, except when simplices contain the origin. Then, we will revert to using Lemma 9.
3 Error bounds for function synthesis
[…]
Additionally, in Section 3.2, an LMI to enforce a scalar inequality constraint on a closed ball about the origin is developed. This result is later used for synthesis of combined CPA and quadratic storage functions in dissipativity analysis.
3.1 LMI Error Bounds
A positive definite error bound matrix is developed for an LMI constraint applied to an $n$-simplex. Enforcing the LMI constraint plus its error bound on the vertex points of an $n$-simplex ($\sigma=\text{co}\{x_{j}\}_{j=0}^{n}$) implies that the LMI holds for all points within that simplex ($x\in\sigma$). While [giesl2012construction] established Lemma 9 as a general error bound for $\mathcal{C}^{2}$ vector-valued functions, it does not translate automatically to an LMI that contains $\mathcal{C}^{2}$ vector-valued functions because the structure of an LMI affects its definiteness. This factor is considered in the following theorem.
Theorem 10.
Consider
[…]
where $x\in\mathbb{R}^{n},$ $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R},$ $\zeta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m},$ $\phi,\zeta\in\mathcal{C}^{2}.$ Let $\sigma\vcentcolon=\text{co}\{x_{j}\}_{j=0}^{n}$ be an n-simplex in $\mathbb{R}^{n}.$ If $x=\sum_{j=0}^{n}\lambda_{j}x_{j}\in\sigma,$ then
$\mathbf{M}(x)$- $\sum_{j=0}^{n}\lambda_{j}\mathbf{M}(x_{j})\preceq$
$\sum_{j=0}^{n}\lambda_{j}\begin{bmatrix}\frac{1}{2}\hat{\phi}_{j}&*&*\\ \mathbf{0}&\frac{1}{2}\Pi^{-1}&*\\ \hat{\zeta}&\mathbf{0}&-2\Pi^{-1}\end{bmatrix}=\vcentcolon\mathbf{E}(x),$ (11)
where $\hat{\phi}_{j}\geq\max_{j\in\mathbb{Z}_{0}^{n}}\left\lvert\overline{\mathbf{H}}_{\phi}(x_{j},x)\right\rvert$, $\hat{\zeta}\in\mathbb{R}^{m\times 1}$ where each element $\hat{\zeta}^{(k)}\geq\max_{j\in\mathbb{Z}_{0}^{n}}\left\lvert\overline{\mathbf{H}}_{\zeta^{(k)}}(x_{j},x)\right\rvert$ for $k\in\mathbb{Z}_{1}^{m}$, and $\Pi\in\mathbb{R}^{m\times m}$ is a diagonal, positive definite matrix with elements $\pi_{k},$ $k\in\mathbb{Z}_{1}^{m}.$
Moreover, if $\mathbf{M}(x_{j})+\mathbf{E}(x_{j})\preceq 0$ holds for all vertex points of $\sigma,$ then $\mathbf{M}(x)\preceq 0$ holds for all $x\in\sigma.$
Theorem 10 bounds the difference between $\mathbf{M}(x)$ at any convex combination of vertex points, $\mathbf{M}(\sum_{i=0}^{n}\lambda_{i}\mathbf{x}_{i})$, and the convex combination of the LMI evaluated at each vertex point, $\sum_{{j}=0}^{n}\lambda_{j}\mathbf{M}(x_{j})$. The proof parallels [giesl2012construction, Proposition 2.2] by developing remainder terms using Taylor’s theorem, but exploits the structure of $\mathbf{M}(x)$ to establish an LMI error bound. Note the negative identity matrix in (10) is essential to enforce negative semi-definiteness.
{pf}
Let $\tilde{\mathbf{E}}(x)=\mathbf{M}(x)-\sum_{j=0}^{n}\!\lambda_{j}\mathbf{M}(x_{j}).$ As in Lemma 17, apply a Theorem 8 to $\phi$ and $\zeta$ about $x$ for each vertex point to get
[…]
which we define as $\tilde{\mathbf{E}}(x).$ If $\tilde{\mathbf{E}}(x)$ was diagonal, bounding $\tilde{\mathbf{E}}(x)$ across a simplex would be as simple as bounding its individual elements. Since that is not the case, the structure of the matrix is leveraged to bound its off-diagonal terms.
Young’s relation [zhou1988robust] is used, with scaling matrix $\bar{\Pi}=\begin{bmatrix}\mathbf{I}&\mathbf{0}\\ \mathbf{0}&\Pi\end{bmatrix}$, to bound the Hermition of $\tilde{\mathbf{E}}(x)$, resulting in
$\tilde{\mathbf{E}}(x)$ $\preceq\frac{1}{2}\sum\limits_{{j}=0}^{n}\lambda_{j}\begin{bmatrix}e_{j}&\mathbf{0}\\ \mathbf{0}&\Pi^{-1}\end{bmatrix}$,
[…]
Now suppose that $\mathbf{M}(x)\preceq 0$ must be imposed on all $x\in\sigma.$ By assumption, $\mathbf{M}(x_{j})+\mathbf{E}(x_{j})\preceq 0$ holds for each vertex of $\sigma$ $(x_{0},\ldots,x_{n}).$ The set of negative semi-definite LMIs is a convex cone [boyd2004convex]. By enforcing $\mathbf{M}(x_{j})+\mathbf{E}(x_{j})\preceq 0$ on each vertex, the expression $\sum_{{j}=0}^{n}\lambda_{j}(\mathbf{M}(x_{j})+\mathbf{E}(x_{j}))\preceq 0$ also holds. The LMI $\mathbf{M}(x)+\mathbf{E}(x)\preceq 0$ implies $\mathbf{M}(x)\preceq 0,$ because $\mathbf{E}(x)\succeq 0$. Therefore, $\mathbf{M}(x)\preceq 0$ for all $x\in\sigma.$
The weighing matrix $\Pi$ in (10) of Theorem 10 allows for a user to have some control over the error bound of the off diagonal terms of $\mathbf{M}(x)$ and can be treated as a design variable in an optimization problem.
In Theorem 10, functions contained within then LMI are linearized about an arbitrary point within a simplex ($x\in\sigma$). However, a similar process as to Lemma 9 can used to linearize the functions about a fixed vertex, $x_{0}.$ The resulting corollary of this process is stated in Section B.
3.2 Quadratic Error Bounds
Section 3.1 bounded matrix behavior across a simplex of a triangulation. Here, the focus shifts to bounding an inequality constraint within a closed two-norm ball about the origin. As in previous sections, application of Taylor’s theorem is used to bound the inequality across the region. However, by focusing on a two-norm ball about the origin, Theorem 11 later expands concepts used in CPA function synthesis to different function types.
Theorem 11.
Consider the inequality
$\zeta(x)^{\top}\zeta(x)+\frac{1}{2}\left(x^{\top}\theta(x)+\theta(x)^{\top}x\right)\leq 0,$ (12)
where $x\in\mathbb{R}^{n}$, $\theta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n},$ $\zeta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m},$ $\theta,\zeta\in\mathcal{C}^{2},$ $\theta(0){=}0,$ and $\zeta(0){=}0.$ Let $\hat{\theta}{\geq}\max_{k\in\mathbb{Z}_{1}^{n},\xi\in B_{\epsilon}(0)}\left\lVert\mathbf{H}_{\theta^{(k)}}(\xi)\right\rVert_{2}$ and $\hat{\zeta}{\geq}\max_{k\in\mathbb{Z}_{1}^{m},\xi\in B_{\epsilon}(0)}\left\lVert\mathbf{H}_{\zeta^{(k)}}(\xi)\right\rVert_{2}$. Let $\mathbf{J}_{\theta}$ and $\mathbf{J}_{\zeta}$ be the Jacobians of $\theta$ and $\zeta$ evaluated at $x=0$. Further, let $\Pi\in\mathbb{R}^{m\times m}$ be a diagonal, positive definite matrix, where $\underline{\pi}$ is the minimal non-zero value of $\Pi^{-1}$. If
$\mathbf{M}_{\epsilon}{=}\begin{bmatrix}\frac{1}{2}(\mathbf{J}_{\theta}\!^{\top}\!\!{+}\mathbf{J}_{\theta}{+}\hat{\theta}\epsilon\mathbf{I})&\mathbf{J}_{\zeta}^{\top}&\hat{\zeta}\epsilon\mathbf{1}_{1\times m}\\ *&-\mathbf{I}+\frac{1}{2}\Pi^{-1}&\mathbf{0}\\ *&*&-2\underline{\pi}\mathbf{I}\end{bmatrix}\preceq 0$ (13)
holds, then (12) holds for all $x\in B_{\epsilon}(0)$.
{pf}
For vector-valued functions, Taylor’s theorem can be applied to each dimension. By applying Taylor’s Theorem about the origin,
$\zeta(x)=$ $\mathbf{J}_{\zeta}x{+}\frac{1}{2}\begin{bmatrix}x^{\top}\mathbf{H}_{\zeta^{(1)}}(z_{\zeta^{(1)}})\\ \vdots\\ x^{\top}\mathbf{H}_{\zeta^{(m)}}(z_{\zeta^{(m)}})\end{bmatrix}x{=\vcentcolon}(\mathbf{A}_{1}{+}\frac{1}{2}\mathbf{A}_{2})x$
and likewise
$\theta(x)=$ $\mathbf{J}_{\theta}x{+}\frac{1}{2}\begin{bmatrix}x^{\top}\mathbf{H}_{\theta^{(1)}}(\mathbf{z}_{\theta^{(1)}})\\ \vdots\\ x^{\top}\mathbf{H}_{\theta^{(n)}}(z_{\theta^{(n)}})\end{bmatrix}x{=\vcentcolon}(\mathbf{B}_{1}{+}\frac{1}{2}\mathbf{B}_{2})x,$
where $\mathbf{A}{=}\mathbf{A}_{1}{+}\frac{1}{2}\mathbf{A}_{2}$, $\mathbf{B}{=}\mathbf{B}_{1}{+}\frac{1}{2}\mathbf{B}_{2}$. Then, (12) is equivalently expressed as $x^{\top}\mathbf{A}^{\top}\mathbf{A}x+\frac{1}{2}(x^{\top}\mathbf{B}x{+}x^{\top}\mathbf{B}^{\top}x)\leq 0$.
By factoring out $x^{\top}$ and $x$ and performing a Schur complement [boyd1994linear] on $\mathbf{A}^{\top}\mathbf{A}$, (12) becomes the LMI
$\begin{bmatrix}\frac{1}{2}(\mathbf{B}{+}\mathbf{B}^{\top})&\mathbf{A}_{1}^{\top}\\ \mathbf{A}_{1}&-\mathbf{I}\end{bmatrix}+\text{He}\left(\begin{bmatrix}\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}\end{bmatrix}^{\top}\begin{bmatrix}\mathbf{0}&\mathbf{0}\\ \frac{1}{2}\mathbf{A}_{2}&\mathbf{0}\end{bmatrix}\right)\preceq 0.$
Both $\mathbf{A}_{2}$ and $\mathbf{B}_{2}$ contain $x$ – meaning that an infinite number of constraints would be needed to enforce the LMI for all $x\in B_{\epsilon}(0).$ Therefore, the constraint can instead be enforced by a singular LMI that bounds it above for all $x\in B_{\epsilon}(0).$
First, Young’s relation, with scaling matrix $\bar{\Pi}=\begin{bmatrix}\pi_{0}&\mathbf{0}\\ \mathbf{0}&\Pi\end{bmatrix}$ ($\pi_{0}>0)$, bounds the Hermition above to produce the LMI
$\begin{bmatrix}\frac{1}{2}(\mathbf{B}{+}\mathbf{B}^{\top}+\mathbf{A}_{2}^{\top}\Pi\mathbf{A}_{2})&\mathbf{A}_{1}^{\top}\\ \mathbf{A}_{1}&-\mathbf{I}+\frac{1}{2}\Pi^{-1}\end{bmatrix}\preceq 0.$
Then, the definition of matrix two norms and the Cauchy Schwarz inequality are used to produce the bounds $\mathbf{A}_{2}^{\top}\Pi\mathbf{A}_{2}\preceq\left\lVert\mathbf{A}_{2}\right\rVert_{2}^{2}\left\lVert\Pi\right\rVert_{2}\mathbf{I}\preceq\hat{\zeta}^{2}\epsilon^{2}\left\lVert\Pi\right\rVert_{2}\mathbf{I}$ and $\mathbf{B}_{2}\preceq\hat{\theta}\epsilon\mathbf{I}$. Finally, perform a Schur complement [boyd1994linear] on $\frac{1}{2}\hat{\zeta}^{2}\epsilon^{2}\left\lVert\Pi\right\rVert_{2}\mathbf{I}$, which results in the term $\frac{2}{\left\lVert\Pi\right\rVert}$ on the diagonal of the matrix. Let $\pi_{M}$ be the maximum term of $\Pi$. For a diagonal, square matrix $\left\lVert\Pi\right\rVert_{2}\leq\pi_{M}$ and therefore, $\frac{2}{\left\lVert\Pi\right\rVert_{2}}\leq-\frac{2}{\pi_{M}}.$ Recall that $\frac{1}{\pi_{M}}$ is the minimum, nonzero term of $\Pi^{-1}$ to get (13). Thus, (13) implies (12). ∎
As in the previous section, it may be desirable in some instances to treat $\Pi$ in (13) as a design variable.
4 Local dissipativity analysis
The error bounds developed in Section 3 are now leveraged to formulate conditions for which a CPA function or combined CPA and quadratic function is the storage function of a dissipative system. These results are then leveraged in Section 4.2 to develop convex optimization problems that simultaneously synthesize a system’s storage function and analyzes common QSR characterizations – such as gain, passivity, and conic bounds.
4.1 Storage function conditions
Theorems 12 and 13 give conditions for a CPA function or a combined CPA and quadratic function, respectively, adhere to the QSR dissipativity condition (5) across a region of the state space. Theorem 13 considers the case where the system dynamics are such that any input affine term disappears at the origin, while Theorem 13 allows for non-zero control affine terms at the origin. This distinction is important, as all error bounds in the optimization problem must be zero at the origin to prevent infeasibility. While this always occurs when the input term disappears, a modified quadratic and CPA storage function (which is quadratic at the origin) is needed for the nonzero case.
Theorem 12.
Consider (1) with $\mathbf{B}=0$ and $\mathbf{D}=0$. Suppose that $f,G,h,J\in\mathcal{C}^{2}$ on the set $\mathcal{X}\in\mathfrak{R}^{n}$. Let $\mathbf{Q}=\mathbf{Q}^{\top}\in\mathbb{R}^{p\times p}$, $\mathbf{S}\in\mathbb{R}^{p\times m}$, and $\mathbf{R}=\mathbf{R}^{\top}\in\mathbb{R}^{m\times m},$ where $\mathbf{Q}\prec 0$ and $\mathbf{R}\succ 0$. Let $\Pi_{1}\in\mathbb{R}^{m\times m}$ and $\Pi_{2},\Pi_{3}\in\mathbb{R}^{p\times p}$ be diagonal, positive definite matrices, where $\underline{\pi}_{z}$ is the minimum diagonal value of $\Pi_{z}^{-1}$. Define $\mathbf{L}=\{l_{i}\}_{i=1}^{m_{\mathcal{T}}}$. Define the CPA function, $V=\{V_{x}\}_{x\in\mathbb{E}_{\mathcal{T}}}$. Consider
$V_{x}$ $\geq 0\quad\forall x\in\mathbb{E}_{x}$ (14a)
$\left\lVert\nabla V_{i}\right\rVert_{1}$ $\leq l_{i}\quad\forall i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}},$ (14b)
$\left\lVert\mathbf{S}\right\rVert_{2}$ $\leq\hat{s}$, (14c)
$\mathbf{M}(x_{i,j})+\overline{\mathbf{E}}(x_{i,j})$ $\preceq 0,\quad\forall i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}},\forall j\in\mathbb{Z}_{0}^{n},$ (14d)
$\quad\quad\quad x_{i,j}\neq 0$
where $\mathbf{M}$ is the QSR dissipativity condition (5) and
[…]
where $\hat{\mathbf{R}}=\frac{1}{2}(\Pi_{1}^{-1}{+}\hat{s}n\sqrt{pm}\theta_{i}c_{i,j}\mathbf{I})$, $\hat{\mathbf{Q}}=\frac{1}{2}(\Pi_{2}^{-1}{+}\Pi_{3}^{-1})$, $\hat{f}_{i,j}=\frac{n}{2}l\beta_{i}c_{i,j}$, $\hat{G}_{i,j}=n\begin{bmatrix}\mu_{1}c_{i,j}&\ldots&\mu_{m}c_{i,j}\end{bmatrix}^{\top}$, $\hat{h}_{i,j}=n\begin{bmatrix}\rho_{1}c_{i,j}&\ldots&\rho_{p}c_{i,j}\end{bmatrix}^{\top}$, and $\hat{J}_{i,j}=n\sqrt{pm}\theta_{i}c_{i,j}\mathbf{I}.$ Further, $c_{i,j}$ is defined by (9) when $0\in\sigma$ and by (33) when $0\notin\sigma,$
$\beta_{i}\geq\max_{p,q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma_{i}}\left\lvert\frac{\partial f^{(p)}}{\partial\mathbf{x}^{(q)}\partial\mathbf{x}^{(r)}}\Big|_{\mathbf{x}=\xi}\right\rvert$, (16)
$\rho_{i,a}\geq\max_{q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma_{i}}\left\lvert\frac{\partial h^{(a)}}{\partial\mathbf{x}^{(q)}\partial\mathbf{x}^{(r)}}\Big|_{\mathbf{x}=\xi}\right\rvert$, (17)
$\mu_{i,k}\geq\max_{p,q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma_{i}}\left\lvert\frac{\partial G_{k}^{(p)}}{\partial\mathbf{x}^{(q)}\partial\mathbf{x}^{(r)}}\Big|_{\mathbf{x}=\xi}\right\rvert$, (18)
and
$\theta_{i}\geq\max_{k\in\mathbb{Z}_{1}^{m},a\in\mathbb{Z}_{1}^{p},q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma_{i}}\left\lvert\frac{\partial J_{k}^{(a)}}{\partial\mathbf{x}^{(q)}\partial\mathbf{x}^{(r)}}\Big|_{\mathbf{x}=\xi}\right\rvert$. (19)
If $V$ satisfies (14), then $V$ satisfies the dissipation HJI (5) for all $x\in\mathcal{X}^{\circ}$ for $\mathbf{Q},$ $\mathbf{S}$, and $\mathbf{R}.$
{pf}
Theorem 10 is applied to the QSR dissipativity condition (5) resulting in Constraint (14d), which ensures that (5) holds over each simplex for $V$. To expand, the (5) is equivalently expressed as
$\mathbf{M}(x)$ $=\begin{bmatrix}\nabla V^{\top}f(x)&\mathbf{0}&\mathbf{0}\\ *&\hat{\mathbf{R}}&\mathbf{0}\\ *&*&\mathbf{Q}^{-1}\end{bmatrix}$
[…]
Consider the simplex $\sigma_{i}$, where $0\notin\sigma_{i}$. Therefore, for any $x\in\sigma_{i},$ the bound on the LMI evaluated within the simplex versus on the vertices, i.e. $\mathbf{M}(x)-\sum_{j=0}^{n}\!\lambda_{j}\mathbf{M}(x_{i,j})$, is
$\mathbf{E}(x)=\begin{bmatrix}\Xi_{1}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\Xi_{2}+\frac{1}{2}\Pi_{1}^{-1}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\frac{1}{2}(\Pi_{2}^{-1}+\Pi_{3}^{-1})\end{bmatrix}$,
where
$\Xi_{1}\geq\frac{1}{2}\sum_{j=0}^{n}\!\lambda_{j}$ $\Bigl(\nabla V^{\top}\overline{\mathbf{H}}_{f}(x_{i,j},x)$
$+\bigl(\frac{1}{2}\nabla V^{\top}\overline{\mathbf{H}}_{G}(x_{i,j},x)-\overline{\mathbf{H}}_{h}(x_{i,j},x)^{\top}\mathbf{S}\bigr)\Pi_{1}$
$\bigl(\frac{1}{2}\overline{\mathbf{H}}_{G}(x_{i,j},x)^{\top}\nabla V-\mathbf{S}\overline{\mathbf{H}}_{h}(x_{i,j},x)\bigr)$
$+\overline{\mathbf{H}}_{h}(x_{i,j},x)^{\top}\Pi_{2}\overline{\mathbf{H}}_{h}(x_{i,j},x)\Bigr)$
and
$\Xi_{2}\succeq\sum_{j=0}^{n}\!\lambda_{j}$ $\Bigl(\mathbf{S}^{\top}\overline{\mathbf{H}}_{J}(x_{i,j},x)$
$+\frac{1}{2}\overline{\mathbf{H}}_{J}(x_{i,j},x)^{\top}\Pi_{3}\overline{\mathbf{H}}_{J}(x_{i,j},x)\Bigr).$
To bound Taylor remainder terms, $\overline{\mathbf{H}}$, used in $\Xi_{1}$ and $\Xi_{2}$, Lemma 17 is used. Care must first be taken to get these functions into the form where the lemmas can be applied.
First, consider $\Xi_{1}.$ Bounding $\frac{1}{2}\sum_{j=0}^{n}\!\lambda_{j}\nabla V^{\top}\overline{\mathbf{H}}_{f}(x_{i,j},x)$ follows the procedure of [giesl2014revised]. By the linearity of $\nabla V$ and Hölder’s inequality, the expression is bounded above by
$\frac{1}{2}\left\lVert\nabla V\right\rVert_{1}\sum_{j=0}^{n}\!\lambda_{j}\left\lVert\overline{\mathbf{H}}_{f}(x_{i,j},x)\right\rVert_{\infty}$. From here, Lemma 17 is be applied. To bound $\frac{1}{2}\sum_{j=0}^{n}\!\lambda_{j}(\frac{1}{2}\nabla V^{\top}\overline{\mathbf{H}}_{G}(x_{i,j},x)-\overline{\mathbf{H}}_{h}(x_{i,j},x)^{\top}\mathbf{S})\Pi_{1}(\frac{1}{2}\overline{\mathbf{H}}_{G}(x_{i,j},x)^{\top}\nabla V-\mathbf{S}\overline{\mathbf{H}}_{h}(x_{i,j},x)),$ apply the Cauchy-Schwarz inequality, the triangle inequality, and the properties of matrix operator norms to get $\frac{1}{2}\left\lVert\Pi\right\rVert_{1}\sum_{j=0}^{n}\!\lambda_{j}\Bigl(\left\lVert\overline{\mathbf{H}}_{h}(x_{i,j},x)\right\rVert_{2}^{2}\left\lVert\mathbf{S}\right\rVert_{2}^{2}+\frac{1}{2}\left\lVert\nabla V^{\top}\overline{\mathbf{H}}_{G}(x_{i,j},x)\right\rVert_{2}^{2}\Bigr).$ Note also that
$\left\lVert\nabla V^{\top}\overline{\mathbf{H}}_{G}(x_{i,j},x)\right\rVert_{2}^{2}\leq\left\lVert\nabla V\right\rVert_{1}^{2}\sum_{k=1}^{m}\left\lVert\overline{\mathbf{H}}_{G_{k}}(x_{i,j},x)\right\rVert_{\infty}^{2}$. From here, Lemma 9 is applied to terms involving $h$ and $G.$ Finally, $\overline{\mathbf{H}}_{h}(x_{i,j},x)^{\top}\Pi_{2}\overline{\mathbf{H}}_{h}(x_{i,j},x)=\sum_{k=1}^{p}\overline{\mathbf{H}}_{h^{(k)}}(x_{i,j},x)^{2}\pi_{2,k}$ can be bounded using the same procedure as in Lemma 17. Therefore,
$\Xi_{1}=$ $\sum_{j=0}^{n}\!\lambda_{j}\Bigl(\hat{f}_{i,j}+\frac{1}{2}\hat{h}_{i,j}^{\top}\Pi_{2}\hat{h}_{i,j}$
$+\frac{1}{2}\left\lVert\Pi_{1}\right\rVert_{2}\hat{s}^{2}\hat{h}_{i,j}^{\top}\hat{h}_{i,j}+\frac{1}{4}\left\lVert\Pi_{1}\right\rVert_{2}l_{i}^{2}\hat{G}_{i,j}^{\top}\hat{G}_{i,j}\Bigl).$
To determine $\Xi_{2},$ use the properties of matrix norms to produce the bound
$\sum_{j=0}^{n}\!\lambda_{j}(\mathbf{S}^{\top}\overline{\mathbf{H}}_{J_{i,j}}+\frac{1}{2}\overline{\mathbf{H}}_{J}(x_{i,j},x)^{\top}\Pi_{3}\overline{\mathbf{H}}_{J}(x_{i,j},x))\preceq$
$\sum_{j=0}^{n}\!\lambda_{j}(\left\lVert\mathbf{S}\right\rVert_{2}\left\lVert\overline{\mathbf{H}}_{J}(x_{i,j},x)\right\rVert_{2}+\frac{1}{2}\left\lVert\overline{\mathbf{H}}_{J}(x_{i,j},x)\right\rVert_{2}^{2}\left\lVert\Pi_{3}\right\rVert_{2})\mathbf{I}.$
Thus,
$\Xi_{2}=$ $\sum_{j=0}^{n}\!\lambda_{j}\Bigl(\hat{s}n\sqrt{pm}\theta_{i}c_{i,j}+\frac{1}{2}\hat{J}_{i,j}^{2}\left\lVert\Pi_{3}\right\rVert_{2}\Bigr)\mathbf{I}.$
Schur complements [boyd1994linear] are performed about
$\frac{1}{2}\left\lVert\Pi_{1}\right\rVert_{2}\hat{s}^{2}\hat{h}_{i,j}^{\top}\hat{h}_{i,j},$ $\frac{1}{4}\left\lVert\Pi_{1}\right\rVert_{2}l_{i}^{2}\hat{G}_{i,j}^{\top}\hat{G}_{i,j}$, $\frac{1}{2}\hat{h}_{i,j}^{\top}\Pi_{2}\hat{h}_{i,j}$, and $\frac{1}{2}\hat{J}_{i,j}^{2}\left\lVert\Pi_{3}\right\rVert_{2}\mathbf{I}$ to get (15), where $\frac{1}{\left\lVert\Pi\right\rVert}$ is equivalent to the minimal diagonal value of $\Pi^{-1}.$
When $x\in\sigma_{i}$, where $0\in\sigma_{i}$, the same procedure can be applied using cor:lmi_bound_gh rather than Theorem 10. Then, bounds on Taylor remainder terms, $\overline{\mathbf{H}}$ are found using Lemma 9 rather than Lemma 17. The only difference between Constraint (14d) when found using this method is that $c_{i,j}$ will be defined using (9) rather than (33).
The values (16), (17), (18), and (19) exist for all simplexes, because $f,G,h,J\in\mathcal{C}^{2}$ and $\sigma\subseteq\mathcal{X}$ is bounded. Furthermore, $c_{i,j}$ is finite on $\sigma,$ because $\mathcal{X}\in\mathfrak{R}^{n}.$ Defining $c_{i,j}$ as in (9) when $0\in\sigma$ ensures that the error bound is 0 at the origin.
By assumption, (14) holds on each $n$-simplex of the $\mathcal{T}$. By Theorem 10, (5) holds for all $\mathbf{x}\in\mathcal{X}^{\circ}$ for some $\mathbf{Q},$ $\mathbf{S}$, and $\mathbf{R}.$
Theorem 13.
Consider (1). Suppose that $f,G,h,J\in\mathcal{C}^{2}$ on the set $\mathcal{X}\in\mathfrak{R}^{n}.$ Let $\mathbf{Q}=\mathbf{Q}^{\top}\in\mathbb{R}^{p\times p}$, $\mathbf{S}\in\mathbb{R}^{p\times m}$, and $\mathbf{R}=\mathbf{R}^{\top}\in\mathbb{R}^{m\times m},$ where $\mathbf{Q}\prec 0$ and $\mathbf{R}\succ 0$. Let $\Pi_{1},\Pi_{4},\Pi_{6}\in\mathbb{R}^{m\times m}$ and $\Pi_{2},\Pi_{3},\Pi_{5},\Pi_{7},\Pi_{8},\Pi_{9}\in\mathbb{R}^{p\times p}$ be diagonal, positive definite matrices, where $\underline{\pi}_{z}$ is the minimum diagonal value of a matrix $\Pi_{z}^{-1}$. Consider the sets $0\in\Psi$ and $B_{\epsilon}(0)$, where $\Psi\subset B_{\epsilon}(0)\subset\mathcal{X}.$ Define a triangulation $\mathcal{T}=\{\sigma_{i}\}_{i=1}^{m_{\mathcal{T}}}$ on $\mathcal{X}\setminus\Psi^{\circ}$. Define a CPA function $V_{C}=\{\bar{V}_{x}\}_{x\in\mathbb{E}_{\mathcal{T}}}$ and a quadratic function $V_{P}=x^{\top}\mathbf{P}x$. Let there exist a function
[…]
Consider
$\bar{V}_{x}$ $\geq x^{\top}\mathbf{P}x\quad\forall x\in\mathbb{E}_{\delta\Psi}$ (21a)
$\bar{V}_{x}$ $\leq x^{\top}\mathbf{P}x\quad\forall x\in\mathbb{E}_{B_{\epsilon}(0)\setminus\Psi}$ (21b)
$\mathbf{P}$ $\succ 0,$ (21c)
$\left\lVert\mathbf{P}\right\rVert_{2}$ $\leq l_{p},$ (21d)
$\left\lVert\mathbf{S}\right\rVert_{2}$ $\leq\hat{s}$, (21e)
$\mathbf{M}_{\epsilon}$ $\preceq 0,$ (21f)
$\bar{V}_{x}$ $\geq 0\quad\forall x\in\mathbb{E}_{\mathcal{T}}$ (21g)
$\left\lVert\nabla\bar{V}_{i}\right\rVert_{1}$ $\leq l_{i}\quad\forall i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}},$ (21h)
$\mathbf{M}(x_{i,j})$ $\preceq 0,\quad\forall i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}},\forall j\in\mathbb{Z}_{0}^{n},$ (21i)
where $\mathbf{L}=\{l_{i}\}_{i=1}^{m_{\mathcal{T}}}\in\mathbb{R},$ $\mathbf{M}$ is defined by (15) and
$\mathbf{M}_{\epsilon}=$ (22)
$\begin{bmatrix}\Xi_{1}&\!\!\!*&*&*&*&*&*&*&*\\ \!\!\!\mathbf{B}^{\top}\mathbf{P}{-}\mathbf{S}^{\top}\mathbf{J}_{h}&\!\!\!\Xi_{2}&*&*&*&*&*&*&*\\ \mathbf{J}_{h}&\!\!\!\!\!\mathbf{D}&\Xi_{3}&*&*&*&*&*&*\\ \bar{\mathbf{J}}_{G}&\!\!\!\mathbf{0}&\mathbf{0}&\!\!\!-\underline{\pi}_{4}\mathbf{I}&*&*&*&*&*\\ \hat{s}\epsilon np^{\frac{1}{2}}\rho_{\epsilon}\mathbf{I}&\!\!\!\mathbf{0}&\mathbf{0}&\mathbf{0}&\!\!\!-2\underline{\pi}_{5}\mathbf{I}&*&*&*&*\\ l_{p}m^{\frac{1}{2}}n^{\frac{3}{2}}\epsilon^{2}\mu_{\epsilon}\mathbf{I}&\!\!\!\mathbf{0}&\mathbf{0}&\mathbf{0}\!\!\!&\mathbf{0}&\!\!\!\!-2\underline{\pi}_{6}\mathbf{I}&*&*&*\\ \epsilon np^{\frac{1}{2}}\rho_{\epsilon}\mathbf{I}&\!\!\!\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\!\!\!\!-2\underline{\pi}_{7}\mathbf{I}&*&*\\ \mathbf{0}&\!\!\!\bar{\mathbf{J}}_{J}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\!\!\!\!-\underline{\pi}_{8}\mathbf{I}&*\\ \mathbf{0}&\!\!\!\bar{\theta}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\!\!\!\!-2\underline{\pi}_{9}\mathbf{I}\end{bmatrix}$. (23)
Here, $\mathbf{J}_{q}$ represents the Jacobian of function $q$ evaluated at 0, $\Xi_{1}=\mathbf{P}\mathbf{J}_{f}+\mathbf{J}_{f}^{\top}\mathbf{P}+l_{p}n^{\frac{3}{2}}\beta_{\epsilon}\epsilon\mathbf{I}_{n}$, $\Xi_{2}=-\mathbf{R}-\mathbf{S}^{\top}\mathbf{D}-\mathbf{D}^{\top}\mathbf{S}+\Pi_{4}^{-1}+\frac{1}{2}(\Pi_{5}^{-1}+\Pi_{6}^{-1})+(2\hat{s}\epsilon\sqrt{\sum_{k=1}^{m}\left\lVert\mathbf{J}_{j_{k}}(0)\right\rVert_{2}^{2}}+\hat{s}(mp)^{\frac{1}{2}}n\epsilon^{2}\theta_{\epsilon})\mathbf{I}$, $\Xi_{3}=\mathbf{Q}^{-1}+\frac{1}{2}(\Pi_{7}^{-1}+\Pi_{9}^{-1})+\Pi_{8}^{-1}$, $\bar{\mathbf{J}}_{G}=l_{p}\epsilon\sqrt{\sum_{k=1}^{m}\left\lVert\mathbf{J}_{g_{k}}(0)\right\rVert_{2}^{2}}\mathbf{I}_{n}$, $\bar{\mathbf{J}}_{J}=\epsilon\sqrt{\sum_{k=1}^{m}\left\lVert\mathbf{J}_{j_{k}}(0)\right\rVert_{2}^{2}}\mathbf{I}$ $\bar{\theta}=m^{\frac{1}{2}}p^{\frac{1}{2}}n\theta_{\epsilon}\epsilon^{2}\mathbf{I}$,
$\beta_{\epsilon}\geq\max_{p,q,r\in\mathbb{Z}_{1}^{n},\xi\in B_{\epsilon}(0)}\left\lvert\frac{\partial^{2}f^{(p)}(\mathbf{x})}{\partial\mathbf{x}^{(q)}\partial\mathbf{x}^{(r)}}\Bigr|_{\mathbf{x}=\xi}\right\rvert$, (24)
$\rho_{\epsilon}\geq\max_{p\in\mathbb{Z}_{1}^{p},q,r\in\mathbb{Z}_{1}^{n},\xi\in B_{\epsilon}(0)}\left\lvert\frac{\partial^{2}h^{(p)}(x)}{\partial x^{(q)}\partial x^{(r)}}\Bigr|_{x=\xi}\right\rvert$, (25)
$\mu_{\epsilon}\geq\max_{k\in\mathbb{Z}_{1}^{m},p,q,r\in\mathbb{Z}_{1}^{n},\xi\in B_{\epsilon}(0)}\left\lvert\frac{\partial^{2}G^{(p)}_{k}(x)}{\partial x^{(q)}\partial x^{(r)}}\Bigr|_{x=\xi}\right\rvert$, (26)
and
$\theta_{\epsilon}\geq\max_{k\in\mathbb{Z}_{1}^{m},p\in\mathbb{Z}_{1}^{p}q,r\in\mathbb{Z}_{1}^{n},\xi\in B_{\epsilon}(0)}\left\lvert\frac{\partial^{2}J^{(p)}_{k}(x)}{\partial x^{(q)}\partial x^{(r)}}\Bigr|_{x=\xi}\right\rvert$. (27)
Further, $\beta_{i},\rho_{i,a}$ and $\mu_{i,k}$ are defined by (16), (17), and (18), and $c_{i,j}$ is defined by (33).
If $V$ satisfies (21), then $V$ satisfies the dissipation HJI (2) for all $x\in\mathcal{X}^{\circ}$ for $\mathbf{Q}$, $\mathbf{S}$, and $\mathbf{R}.$
{pf}
This proof first shows that (2) holds for $V_{C}$ in $\mathcal{X}^{\circ}\setminus\Psi$ and $V_{P}$ in $B_{\epsilon}(0).$ Then, it is shown to hold for $V$ in $\mathcal{X}^{\circ}.$
When $V_{C}, VV$ is a CPA function and the proof follows similarly to Theorem 12 to show that the dissipation HJI holds for $\mathbf{Q},$ $\mathbf{S},$ and $\mathbf{R}$ on $\mathcal{X}\setminus\Psi^{\circ}$ with function $V_{C}$. When $V=V_{P},$ V is quadratic, Theorem 11 must be leveraged to ensure that the dissipation HJI holds on $B_{\epsilon}(0)$.
Constraint (22) creates an upper bound on the QSR dissipativity LMI (5) for all $\mathbf{x}\in B_{\epsilon}(0)$ using Theorem 11. To develop this constraint, perform a Taylor series expansion on each function within the QSR inequality about the origin using $\nabla V=\mathbf{x}^{\top}\mathbf{P}+\mathbf{P}\mathbf{x}$, i.e. $\mathbf{x}^{\top}\mathbf{P}\hat{f}+\hat{f}^{\top}\mathbf{P}\mathbf{x}-\hat{h}^{\top}\mathbf{Q}\hat{h}+(\mathbf{x}^{\top}\mathbf{P}\hat{G}-\hat{h}^{\top}\mathbf{Q}\hat{J}-\hat{h}^{\top}\mathbf{S})\hat{\mathbf{R}}^{-1}(\hat{G}^{\top}\mathbf{P}\mathbf{x}-\hat{J}^{\top}\mathbf{Q}\hat{h}-\mathbf{S}^{\top}\hat{h})\leq 0,$ where $\hat{\mathbf{R}}=\mathbf{R}+\hat{J}^{\top}\mathbf{S}+\mathbf{S}^{\top}\hat{J}+\hat{h}^{\top}\mathbf{Q}\hat{h}$. For example,
$\hat{G}$= $\mathbf{B}+\tilde{G}(\mathbf{x})=\mathbf{B}+\begin{bmatrix}\mathbf{J}_{g_{1}}(0)\mathbf{x}&\ldots\mathbf{J}_{g_{m}}(0)\mathbf{x}\end{bmatrix}$+
$\frac{1}{2}\begin{bmatrix}\begin{bmatrix}\mathbf{x}^{\top}\mathbf{H}_{g^{(1)}_{1}}(\mathbf{z}_{g^{(1)}_{1}})\\ \vdots\\ \mathbf{x}^{\top}\mathbf{H}_{g^{(n)}_{1}}(\mathbf{z}_{g^{(n)}_{1}})\end{bmatrix}\mathbf{x}&\ldots&\begin{bmatrix}\mathbf{x}^{\top}\mathbf{H}_{g^{(1)}_{m}}(\mathbf{z}_{g^{(1)}_{m}})\\ \vdots\\ \mathbf{x}^{\top}\mathbf{H}_{g^{(n)}_{m}}(\mathbf{z}_{g^{(n)}_{m}})\end{bmatrix}\mathbf{x}\end{bmatrix}$,
where the Hessian of each element is evaluated at a convex combination of $\mathbf{x}$ and 0.
Factor $\mathbf{x}$ out on either side of the inequality. Then a Schur complement is performed about terms involving $\hat{\mathbf{R}}$ and another Schur complement is performed about terms involving $\mathbf{Q}.$ Apply the process of Theorem 11. Note that any term on the diagonal of the matrix that contains $\mathbf{x}$ is bounded above using the matrix two norm, e.g. $\mathbf{S}^{\top}\hat{J}\leq\left\lVert\mathbf{S}\right\rVert_{2}\left\lVert\hat{J}\right\rVert_{2}.$ The resulting LMI is
$\begin{bmatrix}\xi_{1}+\mathbf{P}\mathbf{J}_{f}+\mathbf{J}_{f}^{\top}\mathbf{P}&*&*\\ \mathbf{B}^{\top}\mathbf{P}-\mathbf{S}^{\top}\mathbf{J}_{h}&\bar{\mathbf{R}}+\xi_{2}&*\\ \mathbf{J}_{h}&\mathbf{D}&\mathbf{Q}^{-1}+\frac{1}{2}(\Pi_{7}^{-1}+\Pi_{9}^{-1})+\Pi_{8}^{-1}\end{bmatrix}$,
where $\bar{\mathbf{R}}=-\mathbf{R}-\mathbf{S}^{\top}\mathbf{D}-\mathbf{D}^{\top}\mathbf{S}+\Pi_{4}^{-1}+\frac{1}{2}(\Pi_{5}^{-1}+\Pi_{6}^{-1}),$
$\xi_{1}{=}$ $\Bigl(l_{p}^{2}\left\lVert\Pi_{4}\right\rVert_{2}\epsilon^{2}\sum_{k=1}^{m}\left\lVert\mathbf{J}_{g_{k}}(0)\right\rVert_{2}^{2}{+}l_{p}n^{\frac{3}{2}}\beta_{\epsilon}\epsilon$
${+}\frac{1}{2}(\hat{s}^{2}\epsilon^{2}n^{2}p\rho_{\epsilon}^{2}\left\lVert\Pi_{5}\right\rVert_{2}{+}l_{p}^{2}mn^{3}\epsilon^{4}\mu_{\epsilon}^{2}\left\lVert\Pi_{6}\right\rVert_{2}$
${+}\epsilon^{2}n^{2}p\rho_{\epsilon}^{2}\left\lVert\Pi_{7}\right\rVert_{2})\Bigr)\mathbf{I},$
and
$\xi_{2}{=}$ $\Bigl(\left\lVert\Pi_{8}\right\rVert_{2}\epsilon^{2}\sum_{k=1}^{m}\left\lVert\mathbf{J}_{j_{k}}(0)\right\rVert_{2}^{2}+\frac{1}{2}pmn^{2}\theta_{\epsilon}^{2}\epsilon^{4}\left\lVert\Pi_{9}\right\rVert_{2}$
$+2\hat{s}(\epsilon^{2}\sum_{k=1}^{m}\left\lVert\mathbf{J}_{j_{k}}(0)\right\rVert_{2}^{2})^{\frac{1}{2}}+\hat{s}m^{\frac{1}{2}}p^{\frac{1}{2}}n\epsilon^{2}\theta_{\epsilon}\Bigr)\mathbf{I}.$
Perform a Schur complement about each term in $\xi_{1}$, except for the term $l_{p}n^{\frac{3}{2}}\beta_{\epsilon}\epsilon$, and each term in $\xi_{2}$, except $\hat{s}m^{\frac{1}{2}}p^{\frac{1}{2}}n\epsilon^{2}\theta_{\epsilon}\mathbf{I}$, to get (22).
From Theorem 11, (5) holds for all $\mathbf{x}\in B_{\epsilon}(0)$ with $V_{P}$ for $\mathbf{Q},$ $\mathbf{S}$, and $\mathbf{R}.$
Finally, it must be shown that (5) holds for all $x\in\mathcal{X}^{\circ}$ for the function $V.$ The function $V$ is first shown to be continuous. By (21a) and the linearity of $V_{C}$ on each simplex, $V_{C}(x)\geq V_{P}(x)$ for all $x\in\delta\Psi.$ By (21b) and the linearity of $V_{C}$ on each simplex, there is some region, $\Phi\subseteq B_{\epsilon}(0)\setminus\Psi$ and $\Psi\cap\Phi=\emptyset$, where $V_{C}(x)\leq V_{P}(x)$ for all $x\in\Phi.$ By the continuity of $V_{C}$ across $\mathcal{X}\setminus\Psi^{\circ}$ and the intermediate value theorem [fitzpatrick2009advanced, Theorem 3.11], there must exist connected set of points $x\in\Upsilon\subseteq B_{\epsilon}(0)\setminus\Phi$ where $V_{P}(x)=V_{C}(x).$ Therefore, for all $y\in\Upsilon$, $\lim_{x\rightarrow y}V(x)=V(y)$; thus, $V$ is continuous across $\mathcal{X}$ and (5) holds for all $x\in\mathcal{X}^{\circ}$ for for $\mathbf{Q},$ $\mathbf{S}$, and $\mathbf{R}.$
The function, $V$, found in Theorem 12 or Theorem 13 is defined on a subset of the state space. In order for the dynamical system, (1), to be dissipative, it must be shown that the system remains within the region.
Theorem 14.
Consider (1). Let $\Omega\subseteq\mathcal{X}\in\mathfrak{R}^{n}$ be control invariant for the signals $\mathcal{U}_{\Omega}\subseteq\mathcal{L}_{2}^{m}$. If Theorem 12 or Theorem 13 hold over $\mathcal{X},$ then the family of mappings $\mathcal{G}_{\Omega}$ is locally QSR dissipative in $\Omega.$
{pf}
By assumption, the QSR HJI (5) holds for all $x\in\mathcal{X}$ and all $\mathbf{u}\in\mathcal{L}_{2}^{m}.$ Thus, (5) also holds for all $x\in\Omega$ and $\mathbf{u}\in\mathcal{U}_{\Omega}$ and the mappings $\mathcal{G}_{\Omega}(x_{0})$ are locally dissipative for all $x_{0}\in\Omega$.
4.2 Analysis via Optimization
The conditions for a CPA or combined quadratic and CPA function to adhere to the dissipativity LMI (5) developed in Theorems 12 and 13, respectively, provide the foundation to formulate optimization problems for local dissipativity analysis. Problem 4.1 and Problem 4.2 describe a general optimization problem to determine the dissipativity of a system while simultaneously synthesizing a storage function. Problem 4.1 considers systems with no linear control affine terms, while Problem 4.2 considers those with linear control affine terms. In both problems, the generic cost function $J(\alpha,\beta)$ is used, where $\alpha$ and $\beta$ can be determined based of the desired QSR characterization being analyzed. These options are given in Table 1.
Problem 4.1.
Consider the mapping $\mathcal{G}_{\Omega}(x_{0}):\mathcal{U}_{\Omega}\rightarrow\mathcal{L}_{2e}^{p}$ defined by $\mathbf{y}=\mathcal{G}_{\Omega}(x_{0})\mathbf{u}$ and (1). Let $\mathbf{B}=\mathbf{0}$ and $\mathbf{D}=\mathbf{0}$. Define the triangulation $\mathcal{T}=\{\sigma_{i}\}_{i=1}^{m_{\mathcal{T}}}$ over $\Omega\in\mathfrak{R}^{n},$ the CPA function $V=\{V_{\mathbf{x}}\}_{\mathbf{x}\in\mathbb{E}_{\mathcal{T}}}\in\mathbb{R}^{n}$, and the variables $\mathbf{L}=\{l_{i}\}_{i=1}^{m_{\mathcal{T}}}\in\mathbb{R}$, $\alpha\in\mathbb{R},$ and $\beta\in\mathbb{R}.$ Let $\Pi_{1}\in\mathbb{R}^{m\times m}$ and $\Pi_{2},\Pi_{3}\in\mathbb{R}^{p\times p}$ be diagonal matrices. The convex optimization problem is defined as
$\min_{\alpha,\beta,V,\mathbf{L},\Pi}J(\alpha,\beta)$
s.t.
[…]
$\Pi_{i}\succeq 0,\quad\forall i\in\mathbb{Z}_{1}^{3}.$
Problem 4.2.
Consider the mapping $\mathcal{G}_{\Omega}(x_{0}):\mathcal{U}_{\Omega}\rightarrow\mathcal{L}_{2e}^{p}$ defined by $\mathbf{y}=\mathcal{G}_{\Omega}(x_{0})\mathbf{u}$ in (1). Consider the sets $\Omega\in\mathfrak{R}^{n}$ and $\Psi\subset B_{\epsilon}(0)$, where $\epsilon>0.$ Define $\mathcal{T}=\{\sigma_{i}\}_{i=1}^{m_{\mathcal{T}}}$ as the triangulation of $\Omega\setminus\Psi.$ Define $V$ by (20) and the variables $\mathbf{L}=\{l_{i}\}_{i=1}^{m_{\mathcal{T}}}\in\mathbb{R}$, $\alpha\in\mathbb{R},$ and $\beta\in\mathbb{R}.$ Let $\Pi_{1},\Pi_{5},\Pi_{7}\in\mathbb{R}^{m\times m}$ and $\Pi_{2},\Pi_{3},\Pi_{4},\Pi_{6},\Pi_{8},\Pi_{9},\Pi_{10}\in\mathbb{R}^{p\times p}$ be diagonal matrices. The convex optimization problem is defined as
$\min_{\alpha,\beta,V,\mathbf{L}}J(\alpha,\beta)$
s.t.
[…]
$\Pi_{i}\succeq 0,\quad\forall i\in\mathbb{Z}_{1}^{10}.$
IO Characterization QSR Design Variables Objective (0.1)
$\mathcal{L}_{2}$-gain $\mathbf{Q}=-\mathbf{I}$, $\mathbf{S}=\mathbf{0}$, $\mathbf{R}=\gamma^{2}\mathbf{I}$ $\alpha=\gamma^{2}$, $\beta=0$ $J(\alpha,\beta)=\alpha$ $\alpha>0$
Input Strictly Passive $\mathbf{Q}=\mathbf{0}$, $\mathbf{S}=\frac{1}{2}\mathbf{I}$, $\mathbf{R}=-\nu\mathbf{I}$ $\alpha=\nu$, $\beta=0$ $J(\alpha,\beta)=-\alpha$ $\alpha>0$
Output Strictly Passive $\mathbf{Q}=-\rho\mathbf{I}$, $\mathbf{S}=\frac{1}{2}\mathbf{I}$, $\mathbf{R}=\mathbf{0}$ $\alpha=\rho$, $\beta=0$ $J(\alpha,\beta)=-\alpha$ $\alpha>0$
Nondegenerate conic [joshi2002design, bridgeman2014conic] $\mathbf{Q}=-\mathbf{I}$, $\mathbf{S}=\frac{a+b}{2}\mathbf{I},$ $\mathbf{R}=-ab\mathbf{I}$ $\alpha=\frac{a+b}{2},$ $\beta=ab$ $J(\alpha,\beta)=\alpha^{2}-2\beta$ $\beta<0$
Degenerate interior conic [bridgeman2014conic] $\mathbf{Q}=\mathbf{0}$, $\mathbf{S}=\mathbf{I}$, $\mathbf{R}=-d\mathbf{I}$ $\alpha=d$ $J(\alpha,\beta)=-\alpha$ $\alpha<0$
QSR dissipative $\mathbf{Q}\preceq 0,$ any $\mathbf{S}$, $\mathbf{R}\succeq 0$ $\alpha=0$, $\beta=0$ $J(\alpha,\beta)=0$
Table 1: Different QSR dissipativity characteristics can be determined in Problem 4.2 or Problem 4.2 using the design variables, objective function, and additional constraints given in the above table. This updates [hill1977stability, Table 1] to include design criteria for optimization and degenerate conic systems.
5 Solution guarantees
Finally, we establish that we are able to determine a CPA or quadratic storage function for a system that is locally dissipative. It is important to note that the QSR dissipativity holds with some amount of strictness. Theorems 12 and 13 uses error bounds to ensure the validity of a QSR dissipativity characterization. Therefore, the solution found by Problems 1 and 2 will always be conservative at some level.
Theorem 15.
Consider Problem 4.1. Let $\mathcal{G}_{\Omega}$ be locally QSR dissipative on $\Omega$ with $\mathbf{Q}\preceq-\frac{1}{2}(\Pi_{2}+\Pi_{3})^{-1}$ and $\hat{\mathbf{R}}\preceq-\frac{1}{2}(\Pi_{1}^{-1}+\eta\mathbf{I})$, where $\eta\geq\hat{s}n\sqrt{pm}\theta_{i}c_{i,j}+\hat{J}_{i,j}^{2}$ for all $i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}}$ and $j\in\mathbb{Z}_{0}^{n}.$ Let there exist a storage function $\phi\in\mathcal{C}^{2}$ for which (5) holds in $\Omega,$ where $\sup_{\mathbf{x}\in\Omega}\left\lVert\nabla\phi\right\rVert_{2}\leq C$ and $\max_{i,j\in\mathbb{Z}_{1}^{n},\mathbf{x}\in\Omega}\left\lvert\frac{\partial\phi}{\partial\mathbf{x}^{(i)}\partial\mathbf{x}^{(j)}}\Big|_{\mathbf{x}}\right\rvert\leq A$ for some $C,A<\infty.$ Then, Problem 4.1 will find a valid CPA storage function over $\Omega$ for (1) using a finite number of simplices.
{pf}
Let $\bar{t}$ be the maximum length of any simplex in $\mathcal{T}.$ On each simplex $\sigma_{i}$, define
$\bar{\phi}_{i}=\begin{bmatrix}\phi(\mathbf{x}_{i,1})-\phi(\mathbf{x}_{i,0})\\ \vdots\\ \phi(\mathbf{x}_{i,n})-\phi(\mathbf{x}_{i,0})\end{bmatrix}$,
and define $\mathbf{X}_{i}$ as in Lemma 7.
Let each vertex of the function $V$ be defined by the storage function, $\phi$, i.e $V(x_{i,j})=\phi(x_{i,j})$ for all $i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}},$ $j\in\mathbb{Z}_{0}^{n}.$ By Lemma 7, the gradient of the CPA function $\nabla V_{i}$ on each simplex is $\mathbf{X}_{i}^{-1}\bar{\phi}.$
We first aim to bound the distance between $\nabla V_{i}$, a linear approximation of $\nabla\phi$ across $\sigma_{i},$ and the true $\nabla\phi$ evaluated at any $x\in\sigma_{i}$, i.e. $\left\lVert\mathbf{X}_{i}^{-1}\bar{\phi}-\nabla\phi(x)\right\rVert_{2}$. By the properties of matrix operator norms, $\left\lVert\mathbf{X}_{i}^{-1}\bar{\phi}-\nabla\phi(x)\right\rVert_{2}\leq\left\lVert\mathbf{X}_{i}^{-1}\right\rVert_{2}\left\lVert\bar{\phi}-\mathbf{X}_{i}\nabla\phi(x)\right\rVert_{2}.$
First, let $\overline{\mathbf{X}}\geq\left\lVert\mathbf{X}_{i}^{-1}\right\rVert_{2}$ for all $i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}}.$ Recall from step 3 of the proof of [giesl2014revised, Theorem 5], $\left\lVert\mathbf{X}_{i}^{-1}\right\rVert_{2}\leq\frac{1}{\sqrt{\alpha_{i}}},$ where $\alpha_{i}$ is the minimum eigenvalue of $\mathbf{X}_{i}^{\top}\mathbf{X}_{i}$. Thus, $\overline{\mathbf{X}}$ exists and is finite.
To bound $\left\lVert\bar{\phi}-\mathbf{X}\nabla\phi\right\rVert_{2},$ perform Taylor series expansion about $x\in\sigma_{i}$ for each case that $\phi$ is evaluated in $\bar{\phi}$ to get
$\left\lVert\bar{\phi}-\mathbf{X}\nabla\phi\right\rVert_{2}=$ $\frac{1}{2}\left\lVert\begin{bmatrix}\overline{\mathbf{H}}_{\phi}(x_{i,1})-\overline{\mathbf{H}}_{\phi}(x_{i,0})\\ \vdots\\ \overline{\mathbf{H}}_{\phi}(x_{i,n})-\overline{\mathbf{H}}_{\phi}(x_{i,0})\end{bmatrix}\right\rVert_{2}$
$=$ $\frac{1}{2}\sqrt{\sum_{j=1}^{n}\left\lvert\overline{\mathbf{H}}_{\phi}(x_{i,j})-\overline{\mathbf{H}}_{\phi}(x_{i,0})\right\rvert^{2}}.$
By lemma with $\mathbf{H}$ bound, this is bounded above by $An^{\frac{3}{2}}\bar{t}^{2}.$ Thus, $\left\lVert\mathbf{X}_{i}^{-1}\bar{\phi}-\nabla\phi(x)\right\rVert_{2}$
$\leq\left\lVert\mathbf{X}_{i}^{-1}\right\rVert_{2}\left\lVert\bar{\phi}-\mathbf{X}_{i}\nabla\phi(x)\right\rVert_{2}\leq\overline{\mathbf{X}}An^{\frac{3}{2}}\bar{t}^{2}.$
Consider Constraint (14d).For any element of the matrix that contains $\nabla V$, add and subtract $\nabla\phi,$ e.g. $\nabla V^{\top}f(x_{i,j})=(\nabla V-\nabla\phi+\nabla\phi)^{\top}f(x_{i,j}).$ For each simplex, $\nabla V=\mathbf{X}_{i}^{-1}\bar{\phi}.$ Then, by the Schur complement [boyd1994linear], Constraint (14d) is equivalent to the inequality,
$\Lambda+(\mathbf{X}^{-1}\bar{\phi}-\nabla\phi)^{\top}f(x_{i,j})+\frac{1}{2}(\mathbf{X}^{-1}\bar{\phi}-\nabla\phi)^{\top}G(x_{i,j})\tilde{\mathbf{R}}^{-1}$
$\left(G(x_{i,j})^{\top}(\mathbf{X}^{-1}\bar{\phi}-\nabla\phi)-\mathbf{S}^{\top}h(x_{i,j})-J(x_{i,j})^{\top}\tilde{\mathbf{Q}}h(x_{i,j})\right)$
$+\hat{f}_{i,j}+\frac{1}{2\underline{\pi}_{1}}\hat{s}^{2}\hat{h}_{i,j}^{\top}\hat{h}_{i,j}+\frac{1}{4\underline{\pi}_{1}}l_{i}^{2}\hat{G}_{i,j}^{\top}\hat{G}_{i,j}+\frac{1}{2}\hat{h}_{i,j}^{\top}\Pi_{2}h_{i,j}\leq 0,$
where
$\Lambda=$ $\nabla\phi^{\top}f(x_{i,j})-h(x_{i,j})^{\top}\tilde{\mathbf{Q}}h(x_{i,j})$
$+(\frac{1}{2}\nabla\phi^{\top}G(x_{i,j})-h(x_{i,j})^{\top}\mathbf{S}-h(x_{i,j})^{\top}\tilde{\mathbf{Q}}J(x_{i,j}))$
$\tilde{\mathbf{R}}^{-1}(\frac{1}{2}G(x_{i,j})^{\top}\nabla\phi-\mathbf{S}^{\top}h(x_{i,j})-J(x_{i,j})^{\top}\tilde{\mathbf{Q}}h(x_{i,j})),$
which is equivalent to the (5) evaluated at $x_{i,j}$ using the true storage function, $\phi.$ Further, $\tilde{\mathbf{R}}=\hat{\mathbf{R}}+\frac{1}{2}(\Pi_{1}^{-1}+\hat{s}\sqrt{pm}\theta_{i}c_{i,j}\mathbf{I}),$ and $\tilde{\mathbf{Q}}=(\mathbf{Q}^{-1}+\frac{1}{2}\Pi_{2}^{-1}+\frac{1}{2}\Pi_{3}^{-1})^{-1},$ and $\hat{f}_{i,j},$ $\hat{h}_{i,j}$, $\hat{G}_{i,j}$, and $\hat{s}$ are defined by Theorem 12.
Using the properties of matrix operator norms, the inequality is bounded above by
$\Lambda{+}\Bigl(\left\lVert f(x_{i,j})\right\rVert_{2}{+}\left\lVert\mathbf{X}_{i}^{-1}\right\rVert_{2}\left\lVert\bar{\phi}-\mathbf{X}\nabla\phi\right\rVert_{2}\left\lVert\tilde{\mathbf{R}}^{-1}\right\rVert_{2}\left\lVert G(x_{i,j}\right\rVert_{2}$
$+\left\lVert\tilde{\mathbf{R}}^{-1}\right\rVert_{2}\left\lVert h(x_{i,j})\right\rVert_{2}(\left\lVert\mathbf{S}\right\rVert+\left\lVert\tilde{\mathbf{Q}}\right\rVert_{2}\left\lVert J(x_{i,j}\right\rVert_{2})\Bigr)\Bigl(\left\lVert\mathbf{X}_{i}^{-1}\right\rVert_{2}$
$\left\lVert\bar{\phi}-\mathbf{X}\nabla\phi\right\rVert_{2}\Bigr)+\hat{f}_{i,j}+\frac{1}{2\underline{\pi}_{1}}\hat{s}^{2}\hat{h}_{i,j}^{\top}\hat{h}_{i,j}+\frac{1}{4\underline{\pi}_{1}}l_{i}^{2}\hat{G}_{i,j}^{\top}\hat{G}_{i,j}$
$+\frac{1}{2}\hat{h}_{i,j}^{\top}\Pi_{2}\hat{h}_{i,j}.$
The terms $\hat{f}_{i,j},$ $\hat{h}_{i,j},$ and $\hat{G}_{i,j}$ all contain $c_{i,j}$ defined either by (9) if $0\in\sigma_{i}$ or by (33) if $0\notin\sigma.$ By definition, each $c_{i,j}\leq 2\bar{t}^{2}.$ Altogether, the following inequality upper bounds Constraint (14d):
$\Bigl(\left\lVert f(x_{i,j})\right\rVert_{2}{+}\bar{\mathbf{X}}An^{\frac{3}{2}}\bar{t}^{2}\left\lVert\tilde{\mathbf{R}}^{-1}\right\rVert_{2}\left\lVert G(x_{i,j}\right\rVert_{2}$
$+\left\lVert\tilde{\mathbf{R}}^{-1}\right\rVert_{2}\left\lVert h(x_{i,j})\right\rVert_{2}(\left\lVert\mathbf{S}\right\rVert+\left\lVert\tilde{\mathbf{Q}}\right\rVert_{2}\left\lVert J(x_{i,j}\right\rVert_{2})\Bigr)\bar{\mathbf{X}}An^{\frac{3}{2}}\bar{t}^{2}$
$+nl_{i}\beta_{i}\bar{t}^{2}+\frac{2}{\underline{\pi}_{1}}\hat{s}^{2}n^{2}\bar{t}^{4}\sum_{k=1}^{p}\rho_{k}^{2}+\frac{1}{\underline{\pi}_{1}}l_{i}^{2}n^{2}\bar{t}^{4}\sum_{k=1}^{m}\mu_{k}^{2}$
$+2n^{2}\bar{t}^{4}\sum_{k=1}^{p}\rho_{k}^{2}\pi_{2,k}\leq-\Lambda.$
By assumption $\Lambda<0$ at $x\neq 0$ and $\Lambda=0$ at $x=0$, thus $\Lambda\geq 0$ for all $x\in\Omega.$ If the above inequality holds for all $x_{i,j}\neq 0$, then $V$ is a viable storage function. Each term on the left hand side of the inequality contains a term $\bar{t}$, while the term on the right hand side is constant at any $x\in\Omega$. Therefore, $\bar{t}$ approaches 0, the above inequality will hold for any $x_{i,j}\neq 0.$ Therefore, Problem 4.1 will have a solution. ∎
Theorem 16.
Consider Problem 4.2. Let $\mathcal{G}_{\Omega}$ be locally QSR dissipative on $\Omega$ with $\mathbf{Q}\preceq-\frac{1}{2}(\Pi_{7}+\Pi_{8}+2\Pi_{9})^{-1}$ and $\hat{\mathbf{R}}\preceq-\frac{1}{2}(2\Pi_{4}^{-1}+\Pi_{5}^{-1}+\Pi_{6}^{-1})$, where $\eta\geq\hat{s}n\sqrt{pm}\theta_{i}c_{i,j}+\hat{J}_{i,j}^{2}$ for all $i\in\mathbb{Z}_{1}^{m_{\mathcal{T}}}$ and $j\in\mathbb{Z}_{0}^{n}.$ Let there exist a storage function $\Phi(x)=x^{\top}\mathbf{P}x$ $(\mathbf{P}\succ 0)$ such that (5) holds for (1) when linearized about the origin. Then, Problem 4.2 will find a valid storage function, $V_{P}$ over $B_{\epsilon}(0)$ for (1) using a non-zero $\epsilon$.
{pf}
Consider the region $B_{\epsilon}(0).$ In Problem 4.2, Constraint (22) is equivalently expressed as
$x^{\top}$ $\Bigl(\mathbf{J}_{f}^{\top}\mathbf{P}+\mathbf{P}\mathbf{J}_{f}+\omega_{1}-\mathbf{J}_{h}^{\top}\mathbf{Q}^{-1}\mathbf{J}_{h}$
$-(\mathbf{P}\mathbf{B}{-}\mathbf{J}_{h}^{\top}\mathbf{Q}\mathbf{D}{-}\mathbf{J}_{h}^{\top}\mathbf{S})\bar{\mathbf{R}}^{-1}$
$(\mathbf{B}^{\top}\mathbf{P}{-}\mathbf{D}^{\top}\mathbf{Q}\mathbf{J}_{h}{-}\mathbf{S}^{\top}\mathbf{J}_{h})\Bigr)x$
$\leq-x^{\top}\omega_{1}x,$
where $\omega_{1}=(l_{p}^{2}\left\lVert\Pi_{5}\right\rVert_{2}\epsilon^{2}\sum_{k=1}^{m}\left\lVert\mathbf{J}_{g_{k}}(0)\right\rVert_{2}^{2}+l_{p}n^{\frac{3}{2}}\beta_{\epsilon}\epsilon+\frac{1}{2}(\hat{s}^{2}\epsilon^{2}n^{2}p\rho_{\epsilon}^{2}\left\lVert\Pi_{6}\right\rVert_{2}+l_{p}^{2}mn^{3}\epsilon^{4}\mu_{\epsilon}^{2}\left\lVert\Pi_{7}\right\rVert_{2}+\epsilon^{2}n^{2}p\rho_{\epsilon}^{2}\left\lVert\Pi_{8}\right\rVert_{2}))\mathbf{I}$, $\bar{\mathbf{R}}=-\mathbf{R}-\mathbf{S}^{\top}\mathbf{D}-\mathbf{D}^{\top}\mathbf{S}-\mathbf{D}^{\top}\mathbf{Q}\mathbf{D}+\Pi_{5}^{-1}+\frac{1}{2}(\Pi_{6}+\Pi_{7})+(\left\lVert\Pi_{9}\right\rVert_{2}\epsilon^{2}\sum_{k=1}^{m}\left\lVert\mathbf{J}_{j_{k}}(0)\right\rVert_{2}^{2}+\frac{1}{2}pmn^{2}\theta_{\epsilon}^{2}\epsilon^{4}\left\lVert\Pi_{10}\right\rVert_{2}+2\hat{s}\sqrt{\epsilon^{2}\sum_{k=1}^{m}\left\lVert\mathbf{J}_{j_{k}}(0)\right\rVert_{2}^{2}}+\hat{s}\sqrt{mp}n\epsilon^{2}\theta_{\epsilon})\mathbf{I}.$
An upper bound on this constraint can be found when $\bar{\mathbf{R}}$ is replaced with $\hat{\mathbf{R}}=-\mathbf{R}-\mathbf{S}^{\top}\mathbf{D}-\mathbf{D}^{\top}\mathbf{S}-\mathbf{D}^{\top}\mathbf{Q}\mathbf{D}$. Further, if
$x^{\top}$ $\Bigl(\mathbf{J}_{f}^{\top}\mathbf{P}+\mathbf{P}\mathbf{J}_{f}-\mathbf{J}_{h}^{\top}\mathbf{Q}^{-1}\mathbf{J}_{h}$
$+(\mathbf{J}_{h}^{\top}\mathbf{Q}\mathbf{D}+\mathbf{J}_{h}^{\top}\mathbf{S}-\mathbf{P}\mathbf{B})\hat{\mathbf{R}}^{-1}$
$(\mathbf{D}^{\top}\mathbf{Q}\mathbf{J}_{h}+\mathbf{S}^{\top}\mathbf{J}_{h}-\mathbf{B}\mathbf{P})\Bigr)x$
$\leq-m\omega_{1}\epsilon^{2}\leq-x^{\top}\omega_{1}x$
holds, then so does Constraint (22).
Because there is an $\epsilon$ in each term of $\omega_{1}$, there is some $\epsilon$ small enough that the above inequality holds. For this $B_{\epsilon}(0)$, $\mathbf{P}$ will be a viable solution for the quadratic storage function. ∎
6 Numerical examples
We demonstrate the efficacy of our method on three systems. We first find local conic bounds for a one-dimensional nonlinear system and compare to bounds found through ad-hoc ingenuity. Next, we determine the local $\mathcal{L}_{2}$ gain of a two-dimensional pendulum. Finally, we compare our results to that of [summers2013quantitative] for a three-dimensional polynomial system.
6.1 Nonlinear conic system
We consider the nonlinear system
$\mathcal{G}_{\Omega}(x_{0}):\begin{cases}\dot{x}&=k_{1}x^{3}-(k_{1}+k_{2})x+Bu\\ y&=Cx+Du,\end{cases}$ (28)
where $k_{1},k_{2}>0$. Let $\mathcal{X}=[-1,1]$, where $\Omega=[-0.8092,0.8092]$ is an RPI set found using the SOSAnalysis toolbox [packard2009]. When $k_{1}=1$, $k_{2}=2$, $B=1$, $C=1$, and $D=0,$ the system is locally nondegenerate conic, i.e., $\mathcal{G}_{\Omega}\in\text{cone}(a,b)$, where $a=-5.8\times 10^{-9}$ and $b=0.50$ (see Section C). We then implement Problem 4.2 to determine the conic bounds via storage function synthesis for an increasing number of simplices. Specifically, we consider a triangulation over evenly spaced simplices with $\Delta=0.1,0.01,0.001,0.0001$, where $\epsilon=2\Delta.$ As seen in Figure 1, the conic bounds determined by Problem 4.2 closely match the analytical bounds – even with a relatively small number of simplices.
[…]
Figure 1: The analytical upper (b) and lower (a) conic bounds of (28) are determined via Problem 4.2 for an increasing number of simplices. We are able to determine conic bounds of the system that closely match the analytical bounds.
6.2 $\mathcal{L}_{2}$ gain of Pendulum
Consider the dynamics of a pendulum,
$\mathcal{G}_{\Omega}(x_{0}):\begin{cases}\dot{x}_{1}&=x_{2}\\ \dot{x}_{2}&=-\sin x_{1}-x_{2}+u\\ y&=x_{2},\end{cases}$ (29)
where $x\in\mathbb{R}^{n}$, $u\in\mathbb{R}$, and $y\in\mathbb{R}$. Let $\mathcal{X}\supseteq\Omega$ be the triangulated region about the origin, shown in Figure 3, where the RPI set $\Omega$ is found using [lavaei2025L2, Theorem 5], where the system is able to have inputs of $\left\lvert u\right\rvert<0.1942$. Problem 4.2 was applied to system – solving for the $\mathcal{L}_{2}$ gain specifically. Because the pendulum has a sinusoidal term, SOS techniques [summers2013quantitative] cannot be used for analysis. Instead, we compare to the gain found using the global analytical Lyapunov function $V(x)=(1-\cos x_{1})+\frac{1}{2}x_{2}^{2}$, which results in the analytical gain bound $\gamma\geq 1.$ Figure 2 shows the gain found on $\Omega$ for an increasing number of simplices, where the best result is $\gamma=2.221$ Figure 4 shows an example of a storage function found across $\Omega$ with 652 simplices.
[…]
Figure 2: The $\mathcal{L}_{2}$ gain of the pendulum found on a triangulation over $\mathcal{X}\supseteq\Omega$ for an increasing number of simplices. As the triangulation becomes more fine, the gain found by Problem 4.2 becomes closer to the analytical bound.
Figure 3: An example of a triangulation over $\mathcal{X}\supseteq\Omega$, where $\Omega$ is an RPI set denoted by the red line. The set $\Omega$ was found using [lavaei2025L2, Theorem 5] and is valid for inputs where $\left\lvert u\right\rvert<0.1942.$
Figure 4: An example of a storage function over $\mathcal{X}\supseteq\Omega$ found by Problem 4.2.
6.3 Three dimensional polynomial system
Finally, we implemented Problem 4.2 to determine the $\mathcal{L}_{2}$ gain of the following three-dimensional polynomial system,
$\mathcal{G}_{\Omega}(x_{0}):\begin{cases}\dot{x}_{1}&=-x_{1}-x_{3}+x_{2}-x_{3}x_{2}^{2}\\ \dot{x}_{2}&=-x_{2}x_{3}^{2}-x_{2}+u\\ \dot{x}_{3}&=\frac{1}{2}(x_{1}-x_{3})\\ y&=x_{2}\end{cases}$ (30)
over $\mathcal{X}=[-0.5,0.5]\times[-0.5,0.5]\times[-0.5,0.5],$ where $\Omega\subseteq\mathcal{X}$ is a RPI set determined using the SOSAnalysis package [packard2009]. Using 24,576 simplices, the gain was determined to be $\gamma\geq 1.65.$ In contrast, when applying [packard2009], a viable gain was searched for using the SOSAnalysis package using a simple line search. The gain found by [packard2009] was $\gamma\geq 1.01$.
7 Conclusion
This paper presents a general method to determine the local dissipativity properties of nonlinear control affine systems. We develop a convex optimization program that simultaneously optimizes the desired IO characterization of the system and synthesizes the corresponding storage function. Further, we provide guarantees that our analysis method will always be feasible when determining CPA or quadratic storage functions for systems with strict local dissipativity properties.
Appendix A Function Error Bounds
Lemma 17 determines a bound on the behavior of a function across a simplex.
Lemma 17.
Consider $\zeta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m},$ where $\zeta\in\mathcal{C}^{2}$. Let $\sigma\vcentcolon=\text{co}\{\mathbf{x}_{j}\}_{j=0}^{n}$ be an n-simplex in $\mathbb{R}^{n}.$ If $\sum_{j=0}^{n}\lambda_{j}x_{j}\in\sigma$, then
$\left\lVert\zeta(x)-\sum_{j=0}^{n}\lambda_{j}\zeta(x_{j})\right\rVert_{2}$ $=\frac{1}{2}\sum_{j=0}^{n}\!\lambda_{j}\left\lVert\overline{\mathbf{H}}_{\zeta}(x_{j},x)\right\rVert_{2}$ (31)
$\leq\frac{1}{2}n\sum_{j=0}^{n}\!\lambda_{j}c_{j}(\sum_{k=0}^{m}\mu_{k})\text{ and }$
$\left\lVert\zeta(x)-\sum_{j=0}^{n}\lambda_{j}\zeta(x_{j})\right\rVert_{\infty}$ $=\frac{1}{2}\sum_{j=0}^{n}\!\lambda_{j}\left\lVert\overline{\mathbf{H}}_{\zeta}(x_{j},x)\right\rVert_{\infty}$ (32)
$\leq\frac{1}{2}n\max_{k\in\mathbb{Z}_{1}^{m}}\mu_{k}\sum_{j=0}^{n}\!\lambda_{j}c_{j},$
where
$c_{j}=\max_{\nu\in\mathbb{Z}_{0}^{n}}\left\lVert x_{j}-x_{\nu}\right\rVert_{2}^{2}\text{ and }$ (33)
$\mu_{k}\geq\max_{q,r\in\mathbb{Z}_{1}^{n},\xi\in\sigma}\left\lvert\frac{\partial^{2}\zeta^{(k)}(x)}{\partial x^{(q)}\partial x^{(r)}}\Big|_{x=\xi}\right\rvert$. (34)
{pf}
By definition, any point $x\in\sigma$ can be written as a convex combination of the vertices, i.e., $x=\sum_{j=0}^{n}\lambda_{j}x_{j}.$ Let $\tilde{e}(\mathbf{x})=\zeta(x)-\sum_{j=0}^{n}\lambda_{j}\zeta(x_{j}).$ Consider a dimension of $\zeta$, represented as $\zeta^{(k)}$, and apply Theorem 8 about $x$ for each vertex point in $\tilde{e}(\mathbf{x})$ to get
$\tilde{e}(x)=\zeta(x)-\sum_{j=0}^{n}\lambda_{j}$ $\Big(\zeta^{(k)}(x)+\langle\nabla\zeta^{(k)}(x),\Delta\mathbf{x}_{j}\rangle$
$+\frac{1}{2}\overline{\mathbf{H}}_{\zeta^{(k)}}(x_{j},x)\Big).$
In each dimension, the zero-th order term of each function’s Taylor expansion cancel with $\zeta^{(k)}(\mathbf{x})$ and the summed first order terms are zero as $\sum_{j=0}^{n}\lambda_{j}\Delta\mathbf{x}_{j}=\sum_{j=0}^{n}\lambda_{j}\mathbf{x}_{j}-\mathbf{x}=0.$ Altogether,
$\tilde{e}(\mathbf{x})=-\frac{1}{2}\sum_{j=0}^{n}\lambda_{j}\begin{bmatrix}\overline{\mathbf{H}}_{\zeta^{(1)}}(x_{j},x)\\ \vdots\\ \overline{\mathbf{H}}_{\zeta^{(m)}}(x_{j},x)\end{bmatrix}=-\frac{1}{2}\sum_{j=0}^{n}\lambda_{j}\overline{\mathbf{H}}_{\zeta}(x_{j},x).$
To find (31), recall that $\left\lVert\tilde{e}(\mathbf{x})\right\rVert_{2}\leq\left\lVert\tilde{e}(\mathbf{x})\right\rVert_{1}.$ Therefore,
$\left\lVert\tilde{e}(\mathbf{x})\right\rVert_{2}\leq\frac{1}{2}\sum_{j=0}^{n}\lambda_{j}\sum_{k=1}^{m}\left\lvert\overline{\mathbf{H}}_{\zeta^{(k)}}(x_{j},x)\right\rvert$.
Through application of the Cauchy Schwarz inequality and the properties of matrix operator norms,
$\sum_{k=1}^{m}\left\lvert\overline{\mathbf{H}}_{\zeta^{(k)}}(x_{j},x)\right\rvert\leq\sum_{j=0}^{n}\lambda_{j}\sum_{k=1}^{m}\left\lVert\mathbf{H}_{\zeta^{(k)}}(z)\right\rVert_{2}\left\lVert x_{j}-x\right\rVert_{2}^{2}.$
Then, [giesl2012construction, Lemma 2.3], and the bound $\left\lVert x_{j}-x\right\rVert_{2}^{2}\leq\max_{\nu\in\mathbb{Z}_{0}^{n}}\left\lVert\mathbf{x}_{j}-\mathbf{x}_{\nu}\right\rVert_{2}^{2}$ are applied to find (31).
To find (32), instead consider
$\left\lVert\tilde{e}(\mathbf{x})\right\rVert_{\infty}=\frac{1}{2}\sum_{j=0}^{n}\lambda_{j}\max_{k\in\mathbb{Z}_{1}^{m}}\left\lvert\overline{\mathbf{H}}_{\zeta^{(k)}}(x_{j},x)\right\rvert$
The same procedure as above is applied to get (32). ∎
Appendix B LMI Bound
Corollary 18.
Consider (10) as defined in Theorem 10. Let $\sigma\vcentcolon=\text{co}\{x_{j}\}_{j=0}^{n}$ be an n-simplex in $\mathbb{R}^{n}.$ If $x=\sum_{j=0}^{n}\lambda_{j}x_{j}\in\sigma,$ then
$\mathbf{M}(x)$- $\sum_{j=0}^{n}\lambda_{j}\mathbf{M}(x_{j})\preceq$
$\sum_{j=0}^{n}\lambda_{j}\begin{bmatrix}\frac{1}{2}\tilde{\phi}_{j}&*&*\\ \mathbf{0}&\frac{1}{2}\Pi^{-1}&*\\ \tilde{\zeta}&\mathbf{0}&-2\Pi^{-1}\end{bmatrix}=\vcentcolon\mathbf{E}(x),$ (35)
where $\tilde{\phi}_{j}\geq\max_{j\in\mathbb{Z}_{0}^{n}}\left\lvert\overline{\mathbf{H}}_{\phi}(x,x_{0})-\overline{\mathbf{H}}_{\phi}(x_{j},x_{0})\right\rvert$, $\hat{\zeta}\in\mathbb{R}^{m\times 1}$ where each element
$\tilde{\zeta}^{(k)}\geq\max_{j\in\mathbb{Z}_{0}^{n}}\left\lvert\overline{\mathbf{H}}_{\zeta^{(k)}}(x,x_{0})-\overline{\mathbf{H}}_{\zeta^{(k)}}(x_{j},x_{0})\right\rvert$ for $k\in\mathbb{Z}_{1}^{m}$, and $\Pi\in\mathbb{R}^{m\times m}$ is a diagonal, positive definite matrix with elements $\pi_{k},$ $k\in\mathbb{Z}_{1}^{m}.$
Moreover, if $\mathbf{M}(x_{j})+\mathbf{E}(x_{j})\preceq 0$ holds for all vertex points of $\sigma,$ then $\mathbf{M}(x)\preceq 0$ holds for all $x\in\sigma.$
This proof parallels the Taylor series expansion method of [giesl2012construction, Proposition 2.2] for each individual function within $\mathbf{M}(x).$ As with Lemma 9, the bound developed in cor:lmi_bound_gh is essential for enforcing LMIs on simplices that contain the origin as a vertex.
{pf}
Let $\tilde{\mathbf{E}}(x)=\mathbf{M}(x)-\sum_{j=0}^{n}\!\lambda_{j}\mathbf{M}(x_{j}),$ where in $\mathbf{M}(x),$ $\phi(x)=\phi(x_{0})+\nabla\phi(x_{0})^{\top}(x-x_{0})+\frac{1}{2}\overline{\mathbf{H}}_{\phi}(x,x_{0})$ and $\zeta(x)=\zeta(x_{0})+\nabla\zeta(x_{0})^{\top}(x-x_{0})+\frac{1}{2}\overline{\mathbf{H}}_{\zeta}(x,x_{0})$. As in Lemma 17, apply a Theorem 8 to $\phi$ and $\zeta$ about $x$ for each vertex point to get
$\frac{1}{2}\sum\limits_{{j}=0}^{n}\lambda_{j}\!\!\begin{bmatrix}{-}\overline{\mathbf{H}}_{\phi}(x,x_{0}){+}\overline{\mathbf{H}}_{\phi}(x_{j},x_{0})&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}\!\!$
[…]
From here, the proof follows as Theorem 10. ∎
Appendix C Analytical Conic Bound
Lemma 19.
Consider (28) over $\Omega\subseteq\mathcal{X}$, where $\Omega$ is a RPI set and $\mathcal{X}=[-1,1]$. If there exists a $\mathbf{P}>0$ for which
$\begin{bmatrix}-k_{2}\mathbf{P}+C^{\top}C&\frac{1}{2}\mathbf{P}^{\top}B+C^{\top}D-C^{\top}\frac{a+b}{2}\mathbf{I}\\ \frac{1}{2}B^{\top}\mathbf{P}+D^{\top}C-\frac{a+b}{2}\mathbf{I}C&ab\mathbf{I}+D^{\top}D-(a+b)\mathbf{I}D\end{bmatrix}\preceq 0,$ (36)
then $\mathcal{G}_{\Omega}\in\text{cone}(a,b)$.
{pf}
Consider the storage function $V=\frac{1}{2}\mathbf{P}x^{2}.$ For the system to be locally dissipative,
$x^{\top}\mathbf{P}\dot{x}$ $\leq$
$(Cx+Du)^{\top}\mathbf{Q}(Cx+Du)+2(Cx+Du)^{\top}\mathbf{S}u+u^{\top}\mathbf{R}u,$
must be satisfied for all $x\in\Omega$. The above inequality is equivalent to
$k_{1}\mathbf{P}x^{2}-k_{1}\mathbf{P}x^{4}$ $\geq$
$-k_{2}\mathbf{P}x^{2}+x\mathbf{P}Bu$
$-(Cx+Du)^{\top}\mathbf{Q}(Cx+Du)$
$-2(Cx+Du)^{\top}\mathbf{S}u-u^{\top}\mathbf{R}u.$
Because $x^{4}\leq x^{2}$ when $x\in[-1,1],$ the above inequality may be bound above by
$0$ $\geq-(Cx+Du)^{\top}\mathbf{Q}(Cx+Du)-2(Cx+Du)^{\top}\mathbf{S}u$
$-u^{\top}\mathbf{R}u-k_{2}\mathbf{P}x^{2}+x\mathbf{P}Bu.$
Replacing $\mathbf{Q}$, $\mathbf{S}$, and $\mathbf{R}$ with the associated values in Table 1 for non-degenerate interior conic systems and reformulating the above inequality as an LMI results in (36). ∎